### v1

# Autoformalizing and Simulating Game-Theoretic Scenarios using LLM-augmented Agents

Abstract.
Game-theoretic simulations are a versatile tool for exploring interactions of both natural and artificial agents. However, modelling real-world scenarios and developing simulations often require substantial human expertise and effort. To streamline this process, we present a framework that enables the autoformalization of game-theoretic scenarios using agents augmented by large language models (LLMs). In this approach, LLM-augmented agents translate natural language scenario descriptions into executable logic programs that define the rules of each game, validating these programs for syntactic accuracy. A tournament simulation is then conducted, during which the agents test the functionality of the generated games by playing them. When a ground truth payoff matrix is available, an exact semantic validation can also be performed. The validated games can then be used in further simulations to assess the effectiveness of different strategies. We evaluate our approach on a diverse set of 55 natural language descriptions across five well-known $2\times 2$ simultaneous-move games, demonstrating 96% syntactic and 87% semantic correctness in the generated game rules. Additionally, we assess the LLM-augmented agents’ capability to autoformalize strategies for gameplay.
Key words and phrases:
autoformalization, game theory, large language models
1. Introduction
Game-theoretic simulations are a versatile tool for exploring diverse phenomena, ranging from individual human interactions and organisational dynamics to biological processes, robotics, and interactions among artificial agents. However, modelling real-world scenarios and implementing corresponding simulations typically require substantial human expertise and effort. Consider, for instance, the problem of autonomous vehicles coordination: designing a simulation model involves specifying the rules of interaction, the objectives of each agent, the information available to them, and the strategies they may employ, all while accounting for uncertainties and external factors such as traffic laws and pedestrian behaviour; that model subsequently has to be implemented. Recent advancements in Large Language Models (LLMs) enable a new approach to streamline these tasks.
While LLMs have been evaluated as agents in game-theoretic contexts Akata et al. (2023); Fan et al. (2023); Guo (2023); Lorè and Heydari (2023), inherent limitations such as hallucinations Achiam et al. (2023) and logical and arithmetical errors Imani et al. (2023) restrict their utility as direct rational decision-makers. An alternative approach is to leverage LLMs’ capabilities in format translation, including the translation of natural language into formal representations. This process, known as autoformalization Wu et al. (2022), has shown success in creating formal representations for mathematics Wu et al. (2022); He-Yueya et al. (2023); Jiang et al. (2022) and logic Yang et al. (2023); Pan et al. (2023); Cosler et al. (2023); Chen et al. (2023); Feng et al. (2023). In this work, we apply autoformalization to automatically generate formal representations of game-theoretic scenarios.
This work builds on our previous research Mensfelt et al. (2024), where we introduced a basic framework for autoformalizing natural language scenarios that can be modelled using $2\times 2$ simultaneous-move games. In the earlier work, a formal solver was used to validate the syntactic correctness of the generated formal game representations, while semantic correctness was evaluated manually.
In this work, we integrate the previously introduced autoformalization module into a modular agent model and propose a framework for simulating tournaments with these agents. This framework serves two main purposes. First, by running the tournament, we can automatically verify both the functionality of the generated code and its semantic correctness. This verification can be performed approximately, by comparing the agent’s total payoff with the target payoff, or exactly—when the ground truth payoff matrix is available—by checking whether each combination of players’ moves produces the intended payoff. Second, the framework’s modular agent design supports the autoformalization of strategies, enabling the exploration of how different strategies–whether autoformalized or predefined–perform in autoformalized game scenarios. This approach facilitates the evaluation of strategy efficiency across diverse game-theoretic contexts. The main contributions of this work are as follows:
  • An agent model with formal game and strategy representation: We introduce an agent model that incorporates a formal representation of both games and strategies, along with an autoformalization module, developed in the previous work.
  • Three-level agent validation: We provide a mechanism for validating agents for syntactic correctness (through formal solvers), functionality (by employing the code in gameplay), and semantic correctness (by comparing target and achieved payoffs).
  • A simulation framework for autoformalized scenarios: A framework is introduced for simulating autoformalized game scenarios, enabling the investigation and comparison of different strategies – both autoformalized and predefined.
  • Experimental evaluation: We evaluate the framework, demonstrating high syntactic and semantic correctness in the context of $2\times 2$ simultaneous-move games. We also test the capability of the agents to autoformalize 5 strategies.
The remainder of this paper is structured as follows. Sect. 2 introduces the preliminaries, including foundational concepts in game theory, large language models, and general game playing. In Sect. 3 we provide an overview of the proposed framework, followed by Sect. 4, which describes the methods used to evaluate the framework. Sect. 5 presents and analyses the results. Sect. 6 reviews related work in this area. Finally, Sect. 7 summarizes the conclusions and presents directions for future work.
2. Preliminaries
2.1. Game Theory
Game theory is a versatile and popular mathematical modelling framework that captures interaction scenarios pf varying complexity across various disciplines. Rasmusen Rasmusen (2006) composes games using four main components:
  • Players: players consistently and willingly act towards maximising their utility, in full awareness of the situation they are involved in;
  • Actions: mappings between states and outcomes are actions carried out by the player Osborne and Rubinstein (1994);
  • Payoffs: payoffs are quantitative measures over the consequent outcomes of action tuples chosen by players Gibbons (1992);
  • Information: game information describes a player’s knowledge about the available actions, historical moves, and the player’s image of its opponents Osborne and Rubinstein (1994).
We formally define games in this context as the function of a set of $n$ players $N$, a non-empty set of actions $A_{i},\forall i\in N$, and for each player a utility function $u_{i}:A\rightarrow\mathbb{R}$ mapping a player i’s action to the corresponding payoff $\pi_{i}$.
row/col L R
U ($W_{row}$, $W_{col}$) ($X_{row}$, $X_{col}$)
D ($Y_{row}$, $Y_{col}$) ($Z_{row}$, $Z_{col}$)
Table 1. A general normal form game describing the four possible outcomes $W,X,Y,Z$ where the tuples denote the row player’s and column player’s corresponding payoffs, in general: $(\pi_{row},\pi_{col})$
In the case of symmetric games, the payoffs solely depend on the chosen strategies, while with asymmetric games, the outcomes yield different payoffs based on the player’s identity. In our experiments, the Prisoner’s Dilemma, Hawk-Dove, and Stag-Hunt games are symmetric, and we classify actions U and L as “Cooperate” denoted by $C$ and actions D and R as “Defect” denoted by $D$. This allows us to use the well-known terminology from Axelrod’s tournaments to define these games in terms of the four outcomes: $\mathbf{T}:(D,C),\mathbf{R}:(C,C),\mathbf{P}:(D,D),\mathbf{S}:(C,D)$.
  • Prisoner’s Dilemma is characterised by a dominant temptation to defect, represented by the payoff relation $T>R>P>S$,
  • Hawk-Dove entails switched relative payoffs between the punishment for mutual defection and the sucker’s payoff – $T>R>S>P$,
  • Stag Hunt involves a relatively high reward for mutual cooperation over the temptation to defect – $R>T>P>S$.
The Battle of the Sexes and Matching Pennies are asymmetric games. As the outcomes mean different payoffs for row and column, we cannot use the generalization from above and have to define the games from both players’ perspectives:
  • Battle of the Sexes is a coordination game, where both players prefer an agreement but have different preferences over the coordinated outcomes:
    • row: $W>Z>\{X,Y\}$,
    • column: $Z>W>\{X,Y\}$.
  • Matching Pennies: coordinated outcomes reward the row player, while mismatched actions benefit the column player:
    • row: $\{W,Z\}>\{X,Y\}$,
    • column: $\{X,Y\}>\{W,Z\}$.
2.2. Large Language Models
The rapid development of natural language processing (NLP) through the advancements of transformer architectures Gillioz et al. (2020) and pre-trained models Qiu et al. (2020) led to the emergence of Large Language Models. State-of-the-art (SOTA) LLMs are evolving at a swift rate, largely due to the progressive increase in model parameters and training dataset size Zhao et al. (2023). While SOTA LLMs are improving with general use cases, there are approaches that further improve response accuracy and quality. Fine-tuning is the process of training LLMs on a specific custom dataset, increasing the model’s performance in the specific application context captured by the custom data Ziegler et al. (2020). Chain-of-thought prompting guides LLMs to break down their reasoning into simpler steps, often allowing models to deal with more complex requests or specific contexts more accurately Wei et al. (2023).
OpenAI’s GPT-4 Omni OpenAI (2024), employed in this work, is a multi-modal SOTA model, that supports textual, visual and audio data as well. GPT4-o has been shown to perform exceptionally well on reasoning benchmarks Wang and Zhao (2024).
2.3. General Game Playing
General game playing Genesereth et al. (2005) focuses on developing intelligent systems that explicitly represent the rules of arbitrary new games and learn to play them autonomously, without human intervention. The Game Description Language (GDL) has been proposed as a formal, machine-processable language for describing the rules of arbitrary games Love et al. (2006). GDL focused on information games only, so it was extended in GDL-II Thielscher (2010) to cover n-player games with incomplete information and games in extensive normal form Thielscher (2011). GDL-II is based on the standard syntax and semantics of logic programming and characterised by the special keywords shown in Table 2.
role(R) R is a player
init(F) F holds in the initial position
true(F) F holds in the current position
legal(R, M) R can do move M in the current position
does(R, M) player R does move M
next(F) F holds in the next position
terminal the current position is terminal
goal(R, N) R gets N points in the current position
Table 2. Subset of GDL-II (uppercase letters denote variables, lowercase letters denote predicates/function symbols)
A challenge with GDL systems is that learning without human guidance demands complex reasoning. Players must infer the possible actions of others, effectively analysing hypothetical game scenarios before making decisions. Action formalisms like the classical Situation Calculus McCarthy and Hayes (1981) have been developed for precisely this purpose. Formal schemes and inference methods are readily available for Situation Calculus Giacomo et al. (2010); Lesperance et al. (2024), while their deployment in general game playing presupposes a translation from GDL into existing, suitably expressive action languages. One such scheme Schiffel and Thielscher (2011) shows how to fully embed GDL-II into a version of the Situation Calculus based on knowledge fluents Scherl and Levesque (2003). Our game solver, presented later in section 3.2, is inspired by these previous works, to support light GDL specifications that act as target representation for the autoformalization.
3. LLM-augmented Autoformalizing Agents
3.1. Overview
Figure 1. The overview of our agent model. Green dotted lines denote internal communication and solid black lines external communication.
The framework introduced in this work enables the creation of modular autoformalizing agents and supports tournament play using these agents. An overview of the agent model is presented in Fig. 1. A logic programming component implemented in Prolog consists of two modules: a Game module containing game-independent and game-dependent rules, and a Strategy module describing which action to select against an opponent. Together, these modules form a solver (detailed in Section 3.2) that encapsulates the game logic and determines the agent’s next move. A Python wrapper maintains the agent’s interaction with the external environment, manages the game history (tracking each agent’s moves, opponent moves, and payoffs), and updates the solver’s state. This modular design allows for easy substitution of game-dependent rules and strategies, either through predefined configurations or by autoformalizing natural language descriptions. Agents can also be saved to and reloaded from files. The source code and log files generated during the evaluation are available at [^1] https://github.com/dicelab-rhul/autoformalizing-agents.
3.2. Solver
Our solver is based on logic programming, as described in Mensfelt et al. (2024), but repeated here for completeness. It comprises of a game-independent part representing the rules for any extensive-form game, a game-dependent part defining the rules of a particular game using the predicates from the game-independent part, and a set of auxiliary predicates completing the definition of that particular game. We follow standard Prolog conventions to interpret a game: variables are indicated by uppercase letters, while predicates and function symbols by lowercase. The symbol :- is interpreted as if, and the symbol \+ signifies not (negation by failure). An underscore ’_’ represents a variable whose value is ignored within a definition. The game state is then a situation, initially denoted by a constant (e.g. s0). The binary function do(M, S) denotes the situation resulting from the executing move M in situation S, consistent with the Situation Calculus.
3.2.1. Game-independent part
We define all legal transitions of an extended-form game from an initial situation S to a final situation F as follows:
game(F,F):-
final(F).
game(S,F):-
\+ final(S),
legal(M,S),
game(do(M,S),F).
The game terminates when the final game situation F is reached. Otherwise, in a non-final situation S, the game accepts a legal move M, and the game continues in the next do(M,S) situation, until the final situation F is reached. To determine what holds in each legal situation, we use Situation Calculus, represented as:
holds(F, S):-
initially(F, S).
holds(F, do(M, S)):-
effect(F, M, S).
holds(F, do(M, S)):-
holds(F, S),
\+ abnormal(F, M, S).
holds/2 here is similar to true/1 in GDL, but with an additional parameter of the situation in which the fluent is true. It states that a fluent F holds in the initial situation (init/1 in GDL), a new fluent F is initiated by the effects of a move M executed in a situation S (next/1 in GDL), and a fluent F persists after a move is made, provided it is not abnormal; abnormal fluents do not persist (implicit in GDL). We also use rules of the form:
finally(F, S):- Conditions.
to return derived fluents F describing the result of the game, when the Conditions hold in the final situation S.
3.2.2. Game-dependent part
To represent a specific game we need to define game-dependent predicates for the initial state initial/1, the legal moves legal/2, what holds in the initial game situation via initially/2, the effects of a move on a situation via effect/3, what stops persisting in a situation after the execution of a move via abnormal/3, the final situation final/1, and the result of the game via finally/2 definitions. To exemplify these definitions, we show how to describe a PD game to our solver. The initial situation s0 is defined as:
initial(s0).
What holds in this initial situation we specify it as:
initially(player(p1), s0).
initially(player(p2), s0).
initially(role(p1,row), s0).
initially(role(p2,col), s0).
initially(control(p1), s0).
initially(control(p2), s0).
These assertions define first the player names represented by unique identifiers (p1 and p2), their roles (p1 is the row player, while p2 is the column player), and then the fact that initially either of them can the game next (via the control/1 fluent,as in GDL). What holds in the initial situation changes as a result of move being made in the game. We represent moves as terms of the form move(P, M), where P is a player, and M is a move. As in PD it is possible for any player to defect (’D’) or cooperate (’C’), we express this as:
possible(move(P,’D’),S):-holds(player(P),S).
possible(move(P,’C’),S):-holds(player(P),S).
It is then legal for a player to make a possible move if they have the control to execute it in the current situation:
legal(move(P,M), S):-
possible(move(P,M), S),
holds(control(P), S).
As an effect of a legal move M being made by a player P, we record in the next situation that the player did that move (effect/3 is like next/1 in GDL):
effect(did(P, M), move(P, M), S).
Once a legal move is executed, the player loses control, which we specify in our framework as:
abnormal(control(P), move(P, M), S).
When a player loses control cannot play a move until control is passed back. When each player makes a move from the initial situation, the resulting final situation can be determined as:
final(do(M2, do(M1, S))):-
ground(M2),
ground(M1),
initial(S).
Assuming the payoff matrix following the structure presented in Table 1 defined as:
payoff(’D’, D’, 1, 1).
payoff(’C’, D’, 0, 5).
payoff(’D’, C’, 5, 0).
payoff(’C’, C’, 3, 3).
the outcome of the game holds information about the actual moves made by the players, and their payoffs.
finally(outcome(P1,M1,U1,P2,M2,U2), S):-
holds(role(P1, row), S),
holds(did(P1, M1), S),
holds(role(P2, col), S),
holds(did(P2, M2), S),
payoff(M1, M2, U1, U2).
We can then extract specific outcome information, e.g. the player’s utility, through a goal/2 fluent (as in GDL) e.g.:
finally(goal(P1, U1), S):-
finally(outcome(P1,_,U1,_,_,_), S).
finally(goal(P2, U2), S):-
finally(outcome(_,_,_,P2,_,U2), S).
This finalises the definition of PD, enabling us to use the game description above to reason about the game’s dynamics. For instance, if player p1 wanted to analyse the game and identify the actions required to achieve a utility of 5, then it asks the query:
?- game(s0,F), finally(goal(p1,5),F).
Applying the PD payoff matrix, our solver produces two solutions for F, separated by semicolons (;). The result false indicates ”no further answers”:
F = do(move(p2,’C’),do(move(p1,’D’),s0));
F = do(move(p1,’D’),do(move(p2,’C’),s0));
false.
In the first solution, p1 acts first and p2 second, while in the second solution, the order is reversed. This outcome is expected, as both players have initial control in our game definition, resulting in both action order combinations being shown.
3.2.3. Strategy
Strategies are submitted separately from the game-dependent component. In the current implementation, they may rely on specific predicates – default_move/2, opposite_move/2, possible/2, and finally/2 – which are required from the game-dependent solver. The select/4 predicate is used to determine the next move M of player P against opponent O in game situation S, as illustrated in the tit-for-tat strategy example below:
select(P, O, S, M):-
\+ holds(last_move(O, _LMo), S),
holds(default_move(P, M), S).
select(P, O, S, Mo):-
holds(last_move(O, Mo), S).
Here, initially P selects the default move (typically to cooperate), otherwise it mirrors the opponent’s move in the last round. More sophisticated strategies can be expressed. For example, below we show how to express the best response of a player for a game as:
select(P, O, S, M):-
\+ holds(last_move(O, _), S),
holds(default_move(P, M),S).
select(P, O, S, M):-
holds(last_move(O, LMo), S),
findall(Ui-Mi,
(game(S, F),
finally(outcome(P,Mi,Ui,O,Mo,Uo),F),
Ui >= Uo),
Options),
sort(0, @>, Options, Ranked),
highest(Ranked, M).
This second strategy differs from the previous one on tit-for-tat in that its second clause illustrates how to use our representation of games to perform complex reasoning with the game rules. This is done by ranking all legal future outcomes in a given situation according to he utility these have for the player and the opponent, and select the move that gives the highest utility for the player. This strategy also demonstrates the complexity of the autoformalization task.
3.2.4. Constraint checking
We check that the auto-formalized payoff matrix satisfies the structural and content constraints of the specified game (as outlined in Sect. 2.1), by defining validity constraints. The listing below specifies the payoff matrix validity constraints for the Prisoner’s Dilemma.
valid_pd_payoffs(T,R,P,S,C,D):-
payoff(C,C,R,R),
payoff(C,D,S,T),
T>R,
payoff(D,C,T,S),
payoff(D,D,P,P),
R>P,
P>S.
Since the auto-formalization module does not impose constraints on the names of actions, this predicate also identifies which actions correspond to specific roles (e.g., determining that ‘coop’ = C).
3.3. Autoformalization Module
Algorithm 1 Generating game rules predicates from natural language descriptions Mensfelt et al. (2024). PD stands for Prisoner’s Dilemma.
1:Input:
$\Gamma$: game-independent predicates,
$NL_{PD}$: natural language description of PD,
$\xi_{PD}$: game-specific predicates of PD,
$NL_{NG}$: natural language description of a new game.
2:Output:
$\xi_{NG}$: game-specific predicates for the new game.
3:Parameter:
max_attempts: maximum correction attempts.
4:$\text{attempts}\leftarrow 0$
5:$\text{trace}\leftarrow\emptyset$
[…]
7:    $\xi_{NG}\leftarrow\text{LLM.translate}(\Gamma,NL_{PD},\xi_{PD},NL_{NG})$
[…]
9:    if is_valid then
10:         return $\xi_{NG}$
11:    else
[…]
13:         $\xi_{NG}\leftarrow\text{LLM.self\_correct}(\xi_{NG},\text{trace})$
14:    end if
15:    $\text{attempts}\leftarrow\text{attempts}+1$
16:end while
17:return Unable to generate valid predicates within maximum attempts.
To perform autoformalization, a Large Language Model is employed. The framework supports the use of any language model by providing an abstract interface. Autoformalization of a natural language description of a game or strategy is achieved through one-shot prompting, using game-dependent predicates for the Prisoner’s Dilemma and the tit-for-tat strategy as examples. After generating the code, its syntactic correctness is validated. If errors are detected, the LLM receives feedback, including the erroneous lines, error messages, and instructions for correcting the predicates. The max_attempts parameter specifies the maximum number of correction attempts. The overview of autoformalization algorithm is presented in Listing 1.
3.4. Simulation
In our previous work Mensfelt et al. (2024), we manually validated the semantic correctness of generated programs. In this study, we introduce a tournament-based mechanism to automate the semantic correctness validation for scenarios and strategies where a target total payoff or a payoff matrix is available. In this approach, once a scenario is autoformalized, a copy of an agent is created and supplied with a specified strategy. The agent and its copy then play a specified number of rounds (clones mode). Comparing the agent’s total payoff to the target total payoff provides an approximate validation. When a ground truth payoff matrix is available, it enables exact semantic validation by verifying whether the payoff received for each unique combination of moves matches the target payoff for that combination. Additionally, we perform constraint checking to ensure that the payoff matrix satisfies the specified constraints for a specified game.
The tournament can also evaluate which strategy is most successful, on average, within a given game scenario across a specified number of rounds. This is achieved by supplying agent copies with different strategies and having them compete in a round-robin mode, where each agent plays against every other.
The main parameters of a tournament are as follows:
  • game: can be provided by predefined rules or natural language description to autoformalize,
  • strategies: can be provided as a list of predefined strategy rules or a list of natural language descriptions to autoformalize,
  • mode: playing against a single clone or round-robin,
  • agents number,
  • rounds number.
There are two methods for determining the winners of the tournament:
  1. (1)
    Selecting agents which total payoff matches their target payoff – this is applicable if a target payoff is provided for each agent.
  2. (2)
    Selecting agents that achieved the highest total payoff.
Both methods were utilized in the evaluation of the framework, as detailed in the following sections.
4. Methods
4.1. Parameters
Common parameters
LLM GPT-4o
temperature $1$
maximum output tokens $1024$
maximum attempts number $5$
Table 3. Common experimental parameters.
To evaluate the framework, we conducted three experiments: Experiment 1 tested the autoformalization of game scenarios, Experiment 2 employed the autoformalized game rules in tournament’s round-robin mode, and Experiment 3 examined the autoformalization of strategies. Table 3 provides the experimental parameters shared across all experiments, while Table 4 details the parameters unique to each specific experiment.
Parameter Exp. 1 Exp. 2 Exp. 3
games $55$ $5$ $1$
agents per game $5$ $1$ $5$
strategies 1 $6$ $5$
rounds $4$ $10$ $4$
mode clones round-robin clones
Table 4. Experiment-specific parameters.
4.2. Exp. 1: Autoformalization of Game Descriptions
In the first experiment, we used a dataset of 55 natural language game-theoretic scenarios to evaluate the syntactic and semantic accuracy of logic programs generated by agents. The dataset included 5 scenarios that employed common metaphors (e.g., two prisoners in the Prisoner’s Dilemma) for the five canonical games introduced in Section 2.1. The remaining 50 scenarios – 10 for each game – used alternative descriptions that avoided these typical metaphors (e.g., two politicians competing in a campaign rather than two prisoners in the Prisoner’s Dilemma). For each scenario, five agents were generated with autoformalized game rules and then engaged in tournament, using the tit-for-tat strategy against a clone employing the anti-tit-for-tat strategy, testing all four possible pairs of moves in four rounds. Each agent had a maximum of five attempts to produce syntactically correct code; agents failing after five attempts were labelled as syntactically incorrect. Post-tournament, each agent’s individual and total payoffs were compared to the target, with matches indicating semantic correctness.
4.3. Exp. 2: Axelrod’s Tournament
Strategy Description
anti-default-move Always select the move that is the opposite of the default move.
anti-tit-for-tat Start with a default move. Then, select the move that is the opposite of the opponent’s move in the previous round.
best-response Start with a default move. Then, select a move that would give you the highest payoff in response to the opponent’s move in the previous round.
default-move Always select the default move.
random Select one of the possible moves with uniform probability.
tit-for-tat Start with a default move. Then, mirror the opponent’s move in the previous round.
Table 5. Strategies utilised in the evaluation.
In the second experiment, we selected five agents created in Experiment 1, one for each of the five games. For each game, we conducted a tournament based on the design of Axelrod’s tournament Axelrod and Hamilton (1981). In this setup, six copies of each agent were created and assigned one of six strategies; each agent then played against every other agent, including itself. The strategies used are listed in Table 5.
To adapt the strategies to games without explicit cooperation and defection concepts, we defined a “default move” and its “opposite” as the basis for certain strategies. The default moves were: “Cooperate” for the Prisoner’s Dilemma, “Dove” for Hawk-Dove, “Stag” for Stag Hunt, “Opera” for the Battle of the Sexes, and “Heads” for Matching Pennies.
The main metric for evaluating strategy performance was the total payoff achieved by an agent using a given strategy in each game. To allow for cross-game comparisons of strategy effectiveness, total payoffs were normalized with the following formula:
$norm\_payoff=\frac{total\_payoff-min\_total\_payoff}{{max\_total\_payoff-min\_total\_payoff}},$
$Total\_payoff$ refers to the sum of the payoffs an agent collects across all rounds and $min\_total\_payoff$ and $max\_total\_payoff$ represent the minimum and maximum total payoffs achievable, respectively.
4.4. Exp. 3: Autoformalization of Strategies
We then evaluated the semantic correctness of the autoformalized strategies. To achieve this, we used the natural language description of the tit-for-tat strategy, along with its Prolog implementation, as a reference example. For the other five strategies, the autoformalization module was provided with their natural language descriptions, similar to those shown in Table 5.
Each autoformalized strategy was then assigned to an agent with predefined Prisoner’s Dilemma game rules. This agent played four rounds against a clone utilizing the anti-tit-for-tat strategy. The agent’s total payoff was compared to a target payoff specific to each strategy to assess correctness. For the random strategy, correctness was also manually validated. This tournament setup was repeated five times for each strategy to ensure consistency in results.
5. Results and Discussion
5.1. Exp. 1: Autoformalization of Game Descriptions
Attempts 1 2 3 4 5
Count 93% 7% 0 0 12%
Table 6. Distribution of autoformalization attempts number for 275 agents.
Table 6 shows the distribution of attempts needed to generate syntactically correct code for the 275 agents created in the first experiment. For 93% of agents, syntactically correct code was generated on the first attempt. When the correct code was not generated initially, 7% of agents succeeded on the second attempt, while 12% could not produce correct code within the allowed attempts. This latter case highlights a need for further refinement of the feedback loop to improve success rates. However, the examination of the non-aggregated distribution of the invalid agents showed that an agent without syntactically valid game program was generated in no more than 1 in 5 instances.
Figure 2. The percentage of syntactic and semantic correctness of generated game rules for each game.
The overall syntactic correctness rate was 96%, while semantic correctness reached 87%, closely aligning with the results of our previous work (98% and 88%, respectively). Figure 2 depicts the distribution of syntactic and semantic correctness across various games. The Battle of the Sexes and Stag Hunt emerged as the most error-prone scenarios, exhibiting the lowest semantic correctness rates, though still exceeding 80%.
The total-payoff-based approximate validation demonstrated 88% semantic correctness, while the exact individual-payoff-based validation achieved 87%. Five instances were incorrectly labelled as semantically correct by the approximate method, indicating that the results of the approximate and exact methods were closely aligned. A comparison of constraint checking and exact semantic validation revealed a few cases where the autoformalized payoff matrix satisfied the constraints, but errors in other parts of the program – such as role assignment – resulted in incorrect payoff assignments. This demonstrates that exact validation is feasible only when the ground truth payoff matrix is available.
5.2. Exp. 2: Axelrod’s Tournament
Figure 3. Normalized total payoff for each strategy by game.
Figure 3 presents the normalized total payoffs for each strategy across all games, while Table 7 provides the average normalized total payoff for each strategy across games. Overall, the best-response strategy proved to be the most effective on average, followed closely by tit-for-tat. Best-response was most effective in all games except the Battle of the Sexes and the Prisoner’s Dilemma. In the Prisoner’s Dilemma, best-response achieved a lower payoff than the anti-default-move strategy (analogous to “defect”), as it initiated with the “cooperate” move. The results demonstrate that agents with autoformalized rules can effectively simulate and assess the performance of various strategies within different game-theoretic scenarios.
Strategy Payoff [%]
anti-default-move 50.90
anti-tit-for-tat 55.22
best-response 84.08
default-move 73.96
random 61.08
tit-for-tat 79.84
Table 7. Average normalized total payoffs for each strategy.
5.3. Exp. 3: Autoformalization of Strategies
Strategy Correctness [%]
anti-default-move 100.00
anti-tit-for-tat 100.00
best-response 40.00
default-move 100.00
random 60.00
Table 8. Percentage of correctness for each strategy.
Table 8 presents the percentage of correctly autoformalized strategies. The simpler strategies—default-move, anti-default-move, and anti-tit-for-tat—were always correctly autoformalized. In contrast, the random strategy achieved correct autoformalization in 60% of cases, and best-response, which requires selecting the move that maximizes payoff in response to the opponent’s previous move, was correctly autoformalized in 40% of cases. Since all strategies were generated using only the tit-for-tat implementation as an example, increasing the diversity of examples provided to the autoformalization module could potentially improve its accuracy in correctly formalising more complex strategies.
6. Related Work
Originally proposed for agents in game scenarios, the Game Description Language (GDL) offers a promising approach for formalising complex strategic interactions in multi-agent environments by providing a declarative framework to define game rules, legal moves, and outcome conditions Genesereth et al. (2005). Agents that employ GDL are often used to systematically explore game states and strategic options, supporting rigorous analysis across both cooperative and competitive games, as suggested by previous approaches (e.g., see  Schiffel and Thielscher (2009)). However, these approaches typically rely on human-driven formalisation of specific games, a limitation addressed by the LLM-augmented agents and the GDL representation used here, which enable autoformalization and enhance flexibility.
There is a growing body of research exploring the use of LLMs in game theory Lorè and Heydari (2024), systematically examining their potential as player agents in game-theoretic scenarios Fan et al. (2024), utilising their inference abilities to make decisions in games with imperfect information Guo et al. (2023), comparing their behaviour with human participants Fontana et al. (2024), and deploying them in simulation frameworks Mao et al. (2024). While we share with these works an interest in the broad capabilities and value of LLMs in game theoretic settings, our focus diverges by concentrating on how to use LLMs to generate GDL-like executable game specifications. These formal specifications enable agents to rigorously analyse game situations and reason about possible outcomes, providing a structured and robust foundation for decision-making within game-theoretic frameworks. Our work builds on previous research discussed in Mensfelt et al. (2024), significantly extending it by embedding the autoformalization module within an agent model, incorporating the autoformalization of player strategies, implementing a distinct correctness validation process for the autoformalization through tournament play, and conducting experiments to validate the results.
Autoformalization Wu et al. (2022), has shown promising results in translating natural language into mathematical formalisms, enabling the integration of large language models with formal tools for solving reasoning tasksWu et al. (2022); He-Yueya et al. (2023); Jiang et al. (2022). This approach has also been successfully extended to creating logical representations from natural language descriptions. For instance, Yang et al. Yang et al. (2023) applied GPT-3 with few-shot prompting to translate natural language sentences into formal representations, which were then inputted into answer set programs. While effective on NLP benchmarks, this method faced occasional translation errors. To further enhance performance, Pan et al.Pan et al. (2023) combined in-context learning with self-refinement based on solver evaluations, surpassing standard LLM and chain-of-thought prompting approaches on logical reasoning benchmarks. In the context of temporal logic, Cosler et al.Cosler et al. (2023) addressed the problem of translation errors with an interactive system allowing users to refine sub-translations. Fine-tuning LLMs has also proven effective for boosting accuracy in translations to temporal logicsChen et al. (2023) and deductive datasets Feng et al. (2023).
7. Conclusions and Future Work
We have presented a framework for the autoformalization of game-theoretic scenarios, specifically those modelled by $2\times 2$ simultaneous-move games. The framework also enables autoformalization of strategies for these games and the simulation of tournaments based on the autoformalized rules. The generated programs undergo three levels of validation: syntactic validation by consulting a solver, functional testing through gameplay, and semantic validation – either approximate or exact – by comparing achieved payoffs with target outcomes. Evaluation of the framework demonstrated high syntactic and semantic accuracy of the code generated by agents across various interaction scenarios.
Future work will focus on extending the framework to model games beyond $2\times 2$ simultaneous-move scenarios. Our previous work Mensfelt et al. (2024) demonstrated that our sample solver for the Prisoner’s Dilemma could successfully serve to generate code for structurally similar games, such as rock-paper-scissors (with a different number of moves) and sequential versions of the Prisoner’s Dilemma. However, achieving successful generalization to a broader class of games will require assembling a more diverse set of in-context learning examples and developing methods to retrieve the most relevant examples for autoformalization in specific scenarios. The flexibility and generalisability of the introduced framework also make it applicable to games and game-like interactions beyond game theory.
Another direction for future work involves enhancing the feedback loop to more effectively leverage a large language model. At present, the LLM receives feedback solely from the trace generated when consulting the solver with the produced code. In the next phase, we plan to expand this feedback mechanism by incorporating runtime errors, thereby creating a more robust autoformalization module. Finally, in cases where unsupervised autoformalization proves infeasible for complex scenarios, we plan to introduce an interactive mode. This mode will allow users to iteratively refine the generated code alongside the autoformalization module. These interactively generated programs will then be added to our existing program library, expanding the resources available for in-context learning and enabling the autoformalization of increasingly complex scenarios.
{acks}
This work was supported by a Leverhulme Trust International Professorship Grant (LIP-2022-001). The first author would like to thank Bartosz Kosowski for the insightful discussions that contributed to the development of this research.

### v2

@@
-Autoformalizing and Simulating Game-Theoretic Scenarios using LLM-augmented Agents
-Abstract.
-Game-theoretic simulations are a versatile tool for exploring interactions of both natural and artificial agents.
-However, modelling real-world scenarios and developing simulations often require substantial human expertise and effort.
-To streamline this process, we present a framework that enables the autoformalization of game-theoretic scenarios using agents augmented by large language models (LLMs).
-In this approach, LLM-augmented agents translate natural language scenario descriptions into executable logic programs that define the rules of each game, validating these programs for syntactic accuracy.
-A tournament simulation is then conducted, during which the agents test the functionality of the generated games by playing them.
-When a ground truth payoff matrix is available, an exact semantic validation can also be performed.
-The validated games can then be used in further simulations to assess the effectiveness of different strategies.
-We evaluate our approach on a diverse set of 55 natural language descriptions across five well-known $2\times 2$ simultaneous-move games, demonstrating 96% syntactic and 87% semantic correctness in the generated game rules.
-Additionally, we assess the LLM-augmented agents’ capability to autoformalize strategies for gameplay.
-Key words and phrases:
-autoformalization, game theory, large language models
-1. Introduction
-Game-theoretic simulations are a versatile tool for exploring diverse phenomena, ranging from individual human interactions and organisational dynamics to biological processes, robotics, and interactions among artificial agents.
+GAMA: Generative Agents for Multi-Agent Autoformalization
+Abstract
+Multi-agent simulations facilitate the exploration of interactions among both natural and artificial agents.
+However, modelling real-world scenarios and developing simulations often requires substantial expertise and effort.
+To streamline this process, we present a framework that enables the autoformalization of interaction scenarios using agents augmented by large language models (LLMs) utilising game-theoretic formalisms.
+The agents translate natural language descriptions of interactions into executable logic programs that define the rules of each game, ensuring syntactic correctness through validation by a solver.
+A tournament simulation then tests the functionality of the generated game rules and strategies.
+After the tournament, if a ground truth payoff matrix is available, an exact semantic validation is performed.
+We evaluate our approach on a diverse set of 110 natural language descriptions exemplifying five $2\times 2$ simultaneous-move games, achieving 100% syntactic and 76.5% semantic correctness in the generated game rules for Claude 3.5 Sonnet, and 99.82% syntactic and 77% semantic correctness for GPT-4o.
+Additionally, we demonstrate high semantic correctness in autoformalizing gameplay strategies.
+Overall, the results highlight the potential of autoformalization to leverage LLMs in generating formal reasoning modules for decision-making agents.
+1 Introduction
+Multi-agent simulations provide insights into diverse phenomena, ranging from individual human interactions and organizational dynamics to biological processes, robotics, and interactions among artificial agents.
 However, modelling real-world scenarios and implementing corresponding simulations typically require substantial human expertise and effort.
-Consider, for instance, the problem of autonomous vehicles coordination: designing a simulation model involves specifying the rules of interaction, the objectives of each agent, the information available to them, and the strategies they may employ, all while accounting for uncertainties and external factors such as traffic laws and pedestrian behaviour; that model subsequently has to be implemented.
-Recent advancements in Large Language Models (LLMs) enable a new approach to streamline these tasks.
-While LLMs have been evaluated as agents in game-theoretic contexts Akata et al.
+Consider, for instance, the challenge of coordinating autonomous vehicles: creating a simulation model involves specifying the rules of interaction, the goals of each agent, the information available to them, and the strategies they may employ, among other factors.
+Once developed, this model must then be implemented.
+The emergence of Large Language Models (LLMs) enables a novel method for streamlining these tasks through autoformalization.
+We demonstrate it using scenarios that can be modelled as $2\times 2$ simultaneous-move games.
+Although LLMs have been evaluated as agents in game-theoretic contexts Akata et al.
 (2023); Fan et al.
-(2023); Guo (2023); Lorè and Heydari (2023), inherent limitations such as hallucinations Achiam et al.
-(2023) and logical and arithmetical errors Imani et al.
-(2023) restrict their utility as direct rational decision-makers.
-An alternative approach is to leverage LLMs’ capabilities in format translation, including the translation of natural language into formal representations.
+(2023); Guo (2023); Lorè and Heydari (2023), their inherent limitations—such as hallucinations Achiam et al.
+(2023) and logical or arithmetical errors Imani et al.
+(2023)—constrain their effectiveness as rational decision—makers in simulations.
+Rather than relying on LLMs as autonomous agents, a more promising approach is to harness their strengths in format translation, particularly in converting natural language into formal representations.
 This process, known as autoformalization Wu et al.
-(2022), has shown success in creating formal representations for mathematics Wu et al.
+(2022), has shown promise in formalizing mathematics Wu et al.
 (2022); He-Yueya et al.
@@
 (2023).
-In this work, we apply autoformalization to automatically generate formal representations of game-theoretic scenarios.
-This work builds on our previous research Mensfelt et al.
-(2024), where we introduced a basic framework for autoformalizing natural language scenarios that can be modelled using $2\times 2$ simultaneous-move games.
-In the earlier work, a formal solver was used to validate the syntactic correctness of the generated formal game representations, while semantic correctness was evaluated manually.
-In this work, we integrate the previously introduced autoformalization module into a modular agent model and propose a framework for simulating tournaments with these agents.
-This framework serves two main purposes.
-First, by running the tournament, we can automatically verify both the functionality of the generated code and its semantic correctness.
-This verification can be performed approximately, by comparing the agent’s total payoff with the target payoff, or exactly—when the ground truth payoff matrix is available—by checking whether each combination of players’ moves produces the intended payoff.
-Second, the framework’s modular agent design supports the autoformalization of strategies, enabling the exploration of how different strategies–whether autoformalized or predefined–perform in autoformalized game scenarios.
-This approach facilitates the evaluation of strategy efficiency across diverse game-theoretic contexts.
+The resulting formal representations can be utilized in combination with subsymbolic models within a neuro-symbolic framework.
+This approach, as demonstrated by recent advances such as AlphaGeometry, has the potential to achieve human-like problem-solving abilities in challenging domains like geometry Trinh et al.
+(2024).
+In this work, we introduce GAMA: Generative Agents for Multi-Agent Autoformalization, a framework that leverages autoformalization to generate formal representations of strategic interactions.
+These representations serve as the reasoning module of agents deployed in simulations.
+This research builds upon our prior work Mensfelt et al.
+(2024), where we presented a basic framework for autoformalizing natural language interaction scenarios.
+In this paper, we integrate the previously developed autoformalization module into a modular agent architecture within a simulation framework.
 The main contributions of this work are as follows:
-  • An agent model with formal game and strategy representation: We introduce an agent model that incorporates a formal representation of both games and strategies, along with an autoformalization module, developed in the previous work.
-  • Three-level agent validation: We provide a mechanism for validating agents for syntactic correctness (through formal solvers), functionality (by employing the code in gameplay), and semantic correctness (by comparing target and achieved payoffs).
-  • A simulation framework for autoformalized scenarios: A framework is introduced for simulating autoformalized game scenarios, enabling the investigation and comparison of different strategies – both autoformalized and predefined.
-  • Experimental evaluation: We evaluate the framework, demonstrating high syntactic and semantic correctness in the context of $2\times 2$ simultaneous-move games.
-We also test the capability of the agents to autoformalize 5 strategies.
-The remainder of this paper is structured as follows.
-Sect.
-2 introduces the preliminaries, including foundational concepts in game theory, large language models, and general game playing.
-In Sect.
-3 we provide an overview of the proposed framework, followed by Sect.
-4, which describes the methods used to evaluate the framework.
-Sect.
-5 presents and analyses the results.
-Sect.
-6 reviews related work in this area.
-Finally, Sect.
-7 summarizes the conclusions and presents directions for future work.
-2. Preliminaries
-2.1. Game Theory
-Game theory is a versatile and popular mathematical modelling framework that captures interaction scenarios pf varying complexity across various disciplines.
-Rasmusen Rasmusen (2006) composes games using four main components:
+  1. I.
+    An agent model with formal game and strategy representation: We propose an agent model that incorporates formal representations of both games and strategies, along with the autoformalization module from our earlier work.
+Game rules and strategies can either be autoformalized or provided using predefined rules.
+  2. II.
+    Three-level agent validation: We introduce a validation mechanism for agents at three levels: (1) syntactic correctness, using formal solvers; (2) functionality, by testing the code in the simulation; and (3) semantic correctness, by comparing target and achieved payoffs (when a ground truth payoff matrix is available) or verifying compliance with game type constraints (when a ground truth game type is provided).
+  3. III.
+    A simulation framework for autoformalized scenarios: We present a simulation framework for autoformalized interaction scenarios, enabling automated validation and facilitating the exploration and comparison of different strategies for specific scenarios.
+  4. IV.
+    Experimental evaluation: We evaluate the framework, demonstrating high levels of syntactic and semantic correctness for scenarios modelled as $2\times 2$ simultaneous-move games.
+Additionally, we assess the agents’ ability to autoformalize five distinct strategies.
+2 Preliminaries
+2.1 Game Theory
+Game theory is an analytical framework for strategic interactions, where games are composed of four main components Rasmusen (2006):
   • Players: players consistently and willingly act towards maximising their utility, in full awareness of the situation they are involved in;
-  • Actions: mappings between states and outcomes are actions carried out by the player Osborne and Rubinstein (1994);
+  • Actions: mappings between states and outcomes, carried out players Osborne and Rubinstein (1994);
   • Payoffs: payoffs are quantitative measures over the consequent outcomes of action tuples chosen by players Gibbons (1992);
   • Information: game information describes a player’s knowledge about the available actions, historical moves, and the player’s image of its opponents Osborne and Rubinstein (1994).
-We formally define games in this context as the function of a set of $n$ players $N$, a non-empty set of actions $A_{i},\forall i\in N$, and for each player a utility function $u_{i}:A\rightarrow\mathbb{R}$ mapping a player i’s action to the corresponding payoff $\pi_{i}$.
+We formally define games in this context as a tuple $(N,A,\Pi)$ consisting of a set of $n$ players $N$, a non-empty set of actions $A_{i},\forall i\in N$, and for each player $i$ a payoff function $\pi_{i}:\times_{j\in N}A_{j}\rightarrow\mathbb{R}$ mapping action profiles to player payoffs.
 row/col L R
@@
 D ($Y_{row}$, $Y_{col}$) ($Z_{row}$, $Z_{col}$)
-Table 1.
-A general normal form game describing the four possible outcomes $W,X,Y,Z$ where the tuples denote the row player’s and column player’s corresponding payoffs, in general: $(\pi_{row},\pi_{col})$
+Table 1: A general normal form game describing the four possible outcomes $W,X,Y,Z$, where the tuples denote the row player’s and column player’s corresponding payoffs, in general: $(\pi_{row},\pi_{col})$.
 In the case of symmetric games, the payoffs solely depend on the chosen strategies, while with asymmetric games, the outcomes yield different payoffs based on the player’s identity.
-In our experiments, the Prisoner’s Dilemma, Hawk-Dove, and Stag-Hunt games are symmetric, and we classify actions U and L as “Cooperate” denoted by $C$ and actions D and R as “Defect” denoted by $D$.
-This allows us to use the well-known terminology from Axelrod’s tournaments to define these games in terms of the four outcomes: $\mathbf{T}:(D,C),\mathbf{R}:(C,C),\mathbf{P}:(D,D),\mathbf{S}:(C,D)$.
+In our experiments, the Prisoner’s Dilemma, Hawk-Dove, and Stag-Hunt are symmetric, and we classify actions U and L as “Cooperate” denoted by $C$ and actions D and R as “Defect” denoted by $D$.
+This allows us to use the terminology from Axelrod’s tournaments Axelrod and Hamilton (1981) to define these games in terms of the four outcomes: $\mathbf{T}:(D,C),\mathbf{R}:(C,C),\mathbf{P}:(D,D),\mathbf{S}:(C,D)$.
   • Prisoner’s Dilemma is characterised by a dominant temptation to defect, represented by the payoff relation $T>R>P>S$,
@@
     • column: $\{X,Y\}>\{W,Z\}$.
-2.2. Large Language Models
-The rapid development of natural language processing (NLP) through the advancements of transformer architectures Gillioz et al.
+2.2 Large Language Models
+The rapid advancement of natural language processing (NLP), driven by transformer architectures Gillioz et al.
 (2020) and pre-trained models Qiu et al.
-(2020) led to the emergence of Large Language Models.
-State-of-the-art (SOTA) LLMs are evolving at a swift rate, largely due to the progressive increase in model parameters and training dataset size Zhao et al.
+(2020), has led to the emergence of Large Language Models (LLMs).
+State-of-the-art (SOTA) LLMs continue to evolve rapidly, primarily due to the increasing scale of model parameters and training datasets Zhao et al.
 (2023).
-While SOTA LLMs are improving with general use cases, there are approaches that further improve response accuracy and quality.
-Fine-tuning is the process of training LLMs on a specific custom dataset, increasing the model’s performance in the specific application context captured by the custom data Ziegler et al.
-(2020).
-Chain-of-thought prompting guides LLMs to break down their reasoning into simpler steps, often allowing models to deal with more complex requests or specific contexts more accurately Wei et al.
-(2023).
-OpenAI’s GPT-4 Omni OpenAI (2024), employed in this work, is a multi-modal SOTA model, that supports textual, visual and audio data as well.
-GPT4-o has been shown to perform exceptionally well on reasoning benchmarks Wang and Zhao (2024).
-2.3. General Game Playing
+In this work, we employ OpenAI’s GPT-4 Omni OpenAI (2024), a multimodal SOTA model capable of processing text, images, and audio.
+It has demonstrated strong performance on reasoning benchmarks Wang and Zhao (2024).
+Likewise, we utilize Anthropic’s Claude 3.5 Sonnet, another high-performing SOTA model, which excels across multiple benchmarks Anthropic (2024).
+2.3 General Game Playing
 General game playing Genesereth et al.
@@
 GDL focused on information games only, so it was extended in GDL-II Thielscher (2010) to cover n-player games with incomplete information and games in extensive normal form Thielscher (2011).
-GDL-II is based on the standard syntax and semantics of logic programming and characterised by the special keywords shown in Table 2.
-role(R) R is a player
-init(F) F holds in the initial position
-true(F) F holds in the current position
-legal(R, M) R can do move M in the current position
-does(R, M) player R does move M
-next(F) F holds in the next position
-terminal the current position is terminal
-goal(R, N) R gets N points in the current position
-Table 2. Subset of GDL-II (uppercase letters denote variables, lowercase letters denote predicates/function symbols)
 A challenge with GDL systems is that learning without human guidance demands complex reasoning.
@@
 One such scheme Schiffel and Thielscher (2011) shows how to fully embed GDL-II into a version of the Situation Calculus based on knowledge fluents Scherl and Levesque (2003).
-Our game solver, presented later in section 3.2, is inspired by these previous works, to support light GDL specifications that act as target representation for the autoformalization.
-3. LLM-augmented Autoformalizing Agents
-3.1. Overview
-Figure 1. The overview of our agent model. Green dotted lines denote internal communication and solid black lines external communication.
-The framework introduced in this work enables the creation of modular autoformalizing agents and supports tournament play using these agents.
-An overview of the agent model is presented in Fig.
-1.
-A logic programming component implemented in Prolog consists of two modules: a Game module containing game-independent and game-dependent rules, and a Strategy module describing which action to select against an opponent.
-Together, these modules form a solver (detailed in Section 3.2) that encapsulates the game logic and determines the agent’s next move.
-A Python wrapper maintains the agent’s interaction with the external environment, manages the game history (tracking each agent’s moves, opponent moves, and payoffs), and updates the solver’s state.
+Our game solver is inspired by these previous works, to support light GDL specifications that act as target representation for the autoformalization.
+3 Generative Agents for Multi-Agent Autoformalization
+3.1 Overview
+Figure 1: GAMA framework: each agent’s game rules and strategy can be predefined, autoformalized, or loaded from an existing agent.
+The initial agent pool undergoes syntactic validation, and agents that fail to produce syntactically valid code within a defined number of attempts are eliminated (denoted by a crossed circle).
+The remaining agents are validated in the tournament and subsequently undergo semantic validation.
+The GAMA framework introduced in this work enables the creation of modular autoformalizing agents and supports tournament play using these agents (Fig.
+1).
+An agent consists of an autoformalization module and a logic programming component implemented in Prolog, which is divided into two modules: a Game module, containing both game-independent and game-dependent rules, and a Strategy module, which describes the actions to select against an opponent.
+Together, these modules form a solver that encapsulates the game logic and determines the agent’s next move.
+A Python wrapper facilitates the agent’s interaction with the external environment, manages the game history (tracking each agent’s moves, opponent moves, and payoffs), and updates the solver’s state.
 This modular design allows for easy substitution of game-dependent rules and strategies, either through predefined configurations or by autoformalizing natural language descriptions.
-Agents can also be saved to and reloaded from files.
-The source code and log files generated during the evaluation are available at [^1] https://github.com/dicelab-rhul/autoformalizing-agents.
-3.2. Solver
-Our solver is based on logic programming, as described in Mensfelt et al.
-(2024), but repeated here for completeness.
-It comprises of a game-independent part representing the rules for any extensive-form game, a game-dependent part defining the rules of a particular game using the predicates from the game-independent part, and a set of auxiliary predicates completing the definition of that particular game.
-We follow standard Prolog conventions to interpret a game: variables are indicated by uppercase letters, while predicates and function symbols by lowercase.
-The symbol :- is interpreted as if, and the symbol \+ signifies not (negation by failure).
-An underscore ’_’ represents a variable whose value is ignored within a definition.
-The game state is then a situation, initially denoted by a constant (e.g.
-s0).
-The binary function do(M, S) denotes the situation resulting from the executing move M in situation S, consistent with the Situation Calculus.
-3.2.1. Game-independent part
-We define all legal transitions of an extended-form game from an initial situation S to a final situation F as follows:
-game(F,F):-
-final(F).
-game(S,F):-
-\+ final(S),
-legal(M,S),
-game(do(M,S),F).
-The game terminates when the final game situation F is reached.
-Otherwise, in a non-final situation S, the game accepts a legal move M, and the game continues in the next do(M,S) situation, until the final situation F is reached.
-To determine what holds in each legal situation, we use Situation Calculus, represented as:
-holds(F, S):-
-initially(F, S).
-holds(F, do(M, S)):-
-effect(F, M, S).
-holds(F, do(M, S)):-
-holds(F, S),
-\+ abnormal(F, M, S).
-holds/2 here is similar to true/1 in GDL, but with an additional parameter of the situation in which the fluent is true.
-It states that a fluent F holds in the initial situation (init/1 in GDL), a new fluent F is initiated by the effects of a move M executed in a situation S (next/1 in GDL), and a fluent F persists after a move is made, provided it is not abnormal; abnormal fluents do not persist (implicit in GDL).
-We also use rules of the form:
-finally(F, S):- Conditions.
-to return derived fluents F describing the result of the game, when the Conditions hold in the final situation S.
-3.2.2. Game-dependent part
-To represent a specific game we need to define game-dependent predicates for the initial state initial/1, the legal moves legal/2, what holds in the initial game situation via initially/2, the effects of a move on a situation via effect/3, what stops persisting in a situation after the execution of a move via abnormal/3, the final situation final/1, and the result of the game via finally/2 definitions.
-To exemplify these definitions, we show how to describe a PD game to our solver.
-The initial situation s0 is defined as:
-initial(s0).
-What holds in this initial situation we specify it as:
-initially(player(p1), s0).
-initially(player(p2), s0).
-initially(role(p1,row), s0).
-initially(role(p2,col), s0).
-initially(control(p1), s0).
-initially(control(p2), s0).
-These assertions define first the player names represented by unique identifiers (p1 and p2), their roles (p1 is the row player, while p2 is the column player), and then the fact that initially either of them can the game next (via the control/1 fluent,as in GDL).
-What holds in the initial situation changes as a result of move being made in the game.
-We represent moves as terms of the form move(P, M), where P is a player, and M is a move.
-As in PD it is possible for any player to defect (’D’) or cooperate (’C’), we express this as:
-possible(move(P,’D’),S):-holds(player(P),S).
-possible(move(P,’C’),S):-holds(player(P),S).
-It is then legal for a player to make a possible move if they have the control to execute it in the current situation:
-legal(move(P,M), S):-
-possible(move(P,M), S),
-holds(control(P), S).
-As an effect of a legal move M being made by a player P, we record in the next situation that the player did that move (effect/3 is like next/1 in GDL):
-effect(did(P, M), move(P, M), S).
-Once a legal move is executed, the player loses control, which we specify in our framework as:
-abnormal(control(P), move(P, M), S).
-When a player loses control cannot play a move until control is passed back. When each player makes a move from the initial situation, the resulting final situation can be determined as:
-final(do(M2, do(M1, S))):-
-ground(M2),
-ground(M1),
-initial(S).
-Assuming the payoff matrix following the structure presented in Table 1 defined as:
-payoff(’D’, D’, 1, 1).
-payoff(’C’, D’, 0, 5).
-payoff(’D’, C’, 5, 0).
-payoff(’C’, C’, 3, 3).
-the outcome of the game holds information about the actual moves made by the players, and their payoffs.
-finally(outcome(P1,M1,U1,P2,M2,U2), S):-
-holds(role(P1, row), S),
-holds(did(P1, M1), S),
-holds(role(P2, col), S),
-holds(did(P2, M2), S),
-payoff(M1, M2, U1, U2).
-We can then extract specific outcome information, e.g. the player’s utility, through a goal/2 fluent (as in GDL) e.g.:
-finally(goal(P1, U1), S):-
-finally(outcome(P1,_,U1,_,_,_), S).
-finally(goal(P2, U2), S):-
-finally(outcome(_,_,_,P2,_,U2), S).
-This finalises the definition of PD, enabling us to use the game description above to reason about the game’s dynamics.
-For instance, if player p1 wanted to analyse the game and identify the actions required to achieve a utility of 5, then it asks the query:
-?- game(s0,F), finally(goal(p1,5),F).
-Applying the PD payoff matrix, our solver produces two solutions for F, separated by semicolons (;). The result false indicates ”no further answers”:
-F = do(move(p2,’C’),do(move(p1,’D’),s0));
-F = do(move(p1,’D’),do(move(p2,’C’),s0));
-false.
-In the first solution, p1 acts first and p2 second, while in the second solution, the order is reversed.
-This outcome is expected, as both players have initial control in our game definition, resulting in both action order combinations being shown.
-3.2.3. Strategy
+The source code and log files generated during the evaluation are available at [^1] https://github.com/dicelab-rhul/GAMA.
+3.2 Solver
+The solver comprises a game-independent part that represents the rules for any extensive-form game, a game-dependent part that defines the rules of a particular game using the predicates from the game-independent part, and a strategy part to select an action.
+Our solver is implemented in Prolog Körner et al.
+(2022); Wielemaker et al.
+(2012)—a declarative programming language well-suited for high-level problem definitions—but it can easily be substituted with another formal solver.
+The game-independent and game-dependent parts were described in detail in Mensfelt et al.
+(2024).
+To enable the agent incorporating the solver to play, we extended it with the strategy part.
+Additionally, we implemented semantic validation in Prolog using validity constraints.
+The strategy part and constraint-checking mechanism are described below.
+3.2.1 Strategy
 Strategies are submitted separately from the game-dependent component.
-In the current implementation, they may rely on specific predicates – default_move/2, opposite_move/2, possible/2, and finally/2 – which are required from the game-dependent solver.
 The select/4 predicate is used to determine the next move M of player P against opponent O in game situation S, as illustrated in the tit-for-tat strategy example below:
-select(P, O, S, M):-
-\+ holds(last_move(O, _LMo), S),
+select(P, O, S, M) if
+not holds(last_move(O, _LMo), S) and
 holds(default_move(P, M), S).
-select(P, O, S, Mo):-
+select(P, O, S, Mo) if
 holds(last_move(O, Mo), S).
-Here, initially P selects the default move (typically to cooperate), otherwise it mirrors the opponent’s move in the last round.
+Here, a player P selects the default move first (typically to cooperate), otherwise it mirrors the opponent’s move in the last round.
+holds/2 is based on the implementation of the Situation Calculus McCarthy and Hayes (1981).
+The strategy relies on specific predicates – default_move/2 and last_move/2 – which are required from the game-dependent part of the solver.
 More sophisticated strategies can be expressed.
-For example, below we show how to express the best response of a player for a game as:
-select(P, O, S, M):-
-\+ holds(last_move(O, _), S),
-holds(default_move(P, M),S).
-select(P, O, S, M):-
-holds(last_move(O, LMo), S),
-findall(Ui-Mi,
-(game(S, F),
-finally(outcome(P,Mi,Ui,O,Mo,Uo),F),
-Ui >= Uo),
-Options),
-sort(0, @>, Options, Ranked),
-highest(Ranked, M).
-This second strategy differs from the previous one on tit-for-tat in that its second clause illustrates how to use our representation of games to perform complex reasoning with the game rules.
-This is done by ranking all legal future outcomes in a given situation according to he utility these have for the player and the opponent, and select the move that gives the highest utility for the player.
-This strategy also demonstrates the complexity of the autoformalization task.
-3.2.4. Constraint checking
+3.2.2 Validity checking
 We check that the auto-formalized payoff matrix satisfies the structural and content constraints of the specified game (as outlined in Sect.
@@
 The listing below specifies the payoff matrix validity constraints for the Prisoner’s Dilemma.
-valid_pd_payoffs(T,R,P,S,C,D):-
-payoff(C,C,R,R),
-payoff(C,D,S,T),
-T>R,
-payoff(D,C,T,S),
-payoff(D,D,P,P),
-R>P,
+valid_pd_payoffs(T,R,P,S,C,D) if
+payoff(C,C,R,R) and
+payoff(C,D,S,T) and
+T>R and
+payoff(D,C,T,S) and
+payoff(D,D,P,P) and
+R>P and
 P>S.
-Since the auto-formalization module does not impose constraints on the names of actions, this predicate also identifies which actions correspond to specific roles (e.g., determining that ‘coop’ = C).
-3.3. Autoformalization Module
-Algorithm 1 Generating game rules predicates from natural language descriptions Mensfelt et al. (2024). PD stands for Prisoner’s Dilemma.
+Since the autoformalization module does not impose constraints on the names of actions, this predicate also identifies which actions names correspond to specific actions.
+3.3 Autoformalization Module
+Algorithm 1 Generating game rules predicates from natural language descriptions. PD stands for Prisoner’s Dilemma.
 1:Input:
@@
 The framework supports the use of any language model by providing an abstract interface.
-Autoformalization of a natural language description of a game or strategy is achieved through one-shot prompting, using game-dependent predicates for the Prisoner’s Dilemma and the tit-for-tat strategy as examples.
+Autoformalization of a natural language description of a game or strategy is achieved through one-shot prompting, using game-dependent predicates for the Prisoner’s Dilemma or the tit-for-tat strategy as examples, respectively.
 After generating the code, its syntactic correctness is validated.
@@
 The max_attempts parameter specifies the maximum number of correction attempts.
-The overview of autoformalization algorithm is presented in Listing 1.
-3.4. Simulation
-In our previous work Mensfelt et al.
-(2024), we manually validated the semantic correctness of generated programs.
-In this study, we introduce a tournament-based mechanism to automate the semantic correctness validation for scenarios and strategies where a target total payoff or a payoff matrix is available.
-In this approach, once a scenario is autoformalized, a copy of an agent is created and supplied with a specified strategy.
-The agent and its copy then play a specified number of rounds (clones mode).
-Comparing the agent’s total payoff to the target total payoff provides an approximate validation.
-When a ground truth payoff matrix is available, it enables exact semantic validation by verifying whether the payoff received for each unique combination of moves matches the target payoff for that combination.
-Additionally, we perform constraint checking to ensure that the payoff matrix satisfies the specified constraints for a specified game.
-The tournament can also evaluate which strategy is most successful, on average, within a given game scenario across a specified number of rounds.
-This is achieved by supplying agent copies with different strategies and having them compete in a round-robin mode, where each agent plays against every other.
-The main parameters of a tournament are as follows:
-  • game: can be provided by predefined rules or natural language description to autoformalize,
-  • strategies: can be provided as a list of predefined strategy rules or a list of natural language descriptions to autoformalize,
-  • mode: playing against a single clone or round-robin,
-  • agents number,
-  • rounds number.
-There are two methods for determining the winners of the tournament:
-  1. (1)
-    Selecting agents which total payoff matches their target payoff – this is applicable if a target payoff is provided for each agent.
-  2. (2)
-    Selecting agents that achieved the highest total payoff.
-Both methods were utilized in the evaluation of the framework, as detailed in the following sections.
-4. Methods
-4.1. Parameters
-Common parameters
-LLM GPT-4o
+The overview of autoformalization algorithm is presented in Listing 1 Mensfelt et al.
+(2024).
+3.4 Simulation
+The framework is employed in this work in two primary modes: validation and simulation.
+These modes facilitate the evaluation of both the semantic correctness of generated programs and the effectiveness of strategies within various game scenarios.
+In validation mode, the framework ensures the semantic correctness of generated programs by comparing their performance against predefined criteria.
+Once a scenario is autoformalized, an agent is created and provided with a specified strategy.
+The agent then plays against a clone of itself for a set number of rounds.
+If a ground truth payoff matrix is available, the framework performs semantic validation by verifying that the payoff for each unique combination of moves matches the target payoff for that combination.
+Additionally, if the ground truth type of the game is known, the framework enforces constraint checking to confirm that the payoff matrix adheres to all specified constraints for the given game.
+In simulation mode, the framework evaluates and compares strategies to determine their effectiveness within a game scenario.
+Agents equipped with different strategies compete in a round-robin tournament, where each agent plays against every other.
+By simulating multiple rounds, the framework identifies the strategy that performs best on average, offering insights into the relative strengths and weaknesses of competing approaches.
+The parameters of the tournament are as follows:
+  • Agent Pool: Contains all agents participating in the tournament. Each agent can be equipped with game rules and strategies either through autoformalization or predefined configurations.
+  • Number of Rounds: Specifies how many rounds each agent will play.
+  • Match Maker: A function that determines the pairing of agents for matches.
+  • Target Payoffs: Optional parameter used for semantic validation.
+The framework provides two methods for determining the tournament’s outcome:
+  1. 1.
+    Selecting agents whose total payoff matches their target payoff, applicable when target payoffs are specified (validation).
+  2. 2.
+    Selecting agents that achieve the highest total payoff, suitable for assessing the success of strategies (simulation).
+Furthermore, the customizable Match Maker function allows for experiments with tailored matching protocols
+4 Methods
+4.1 Parameters
+Parameter Value
+LLM Claude 3.5 Sonnet
+GPT-4o
 temperature $1$
-maximum output tokens $1024$
+maximum output tokens $2048$
 maximum attempts number $5$
-Table 3. Common experimental parameters.
-To evaluate the framework, we conducted three experiments: Experiment 1 tested the autoformalization of game scenarios, Experiment 2 employed the autoformalized game rules in tournament’s round-robin mode, and Experiment 3 examined the autoformalization of strategies.
+Table 2: Common experimental parameters.
+Parameter Exp. 1 Exp. 2 Exp. 3 Exp. 4
+games $55$ $55$ $5$ $5$
+agents in game $5$ $5$ $1$ $1$
+strategies $1$ $1$ $6$ $6$
+rounds $4$ $4$ $10$ $10$
+match maker S S RR S
+Table 3: Experiment-specific parameters. S stands for a mode in which an agent plays against its clone with an opposite strategy, while RR stands for round robin.
+To evaluate the framework, we conducted four experiments: i) autoformalization of game scenarios with numerical (Exp.
+1) and ii) non-numerical (Exp.
+2) payoffs, iii) using the autoformalized game rules in a round-robin tournament (Exp.3), and iv) autoformalization of strategies (Exp.
+4).
 Table 2 provides the experimental parameters shared across all experiments, while Table 3 details the parameters unique to each specific experiment.
-Parameter Exp. 1 Exp. 2 Exp. 3
-games $55$ $5$ $1$
-agents per game $5$ $1$ $5$
-strategies 1 $6$ $5$
-rounds $4$ $10$ $4$
-mode clones round-robin clones
-Table 4. Experiment-specific parameters.
-4.2. Exp. 1: Autoformalization of Game Descriptions
-In the first experiment, we used a dataset of 55 natural language game-theoretic scenarios to evaluate the syntactic and semantic accuracy of logic programs generated by agents.
+4.2 Autoformalization of Interaction Descriptions (Numerical Payoffs)
+We used a dataset of 55 natural language game-theoretic scenarios with numerical payoffs to evaluate the syntactic and semantic accuracy of generated logic programs.
 The dataset included 5 scenarios that employed common metaphors (e.g., two prisoners in the Prisoner’s Dilemma) for the five canonical games introduced in Section 2.1.
-The remaining 50 scenarios – 10 for each game – used alternative descriptions that avoided these typical metaphors (e.g., two politicians competing in a campaign rather than two prisoners in the Prisoner’s Dilemma).
-For each scenario, five agents were generated with autoformalized game rules and then engaged in tournament, using the tit-for-tat strategy against a clone employing the anti-tit-for-tat strategy, testing all four possible pairs of moves in four rounds.
+The remaining 50 scenarios—10 for each game—used alternative descriptions that avoided these typical metaphors (e.g., two politicians competing in a campaign rather than two prisoners in the Prisoner’s Dilemma).
+For each scenario, five agents autoformalized game rules and then played tournament against their clones, generating all four possible pairs of moves in four rounds.
 Each agent had a maximum of five attempts to produce syntactically correct code; agents failing after five attempts were labelled as syntactically incorrect.
 Post-tournament, each agent’s individual and total payoffs were compared to the target, with matches indicating semantic correctness.
-4.3. Exp. 2: Axelrod’s Tournament
+4.3 Autoformalization of Interaction Descriptions (Non-numerical Payoffs)
+This experiment was analogous to the previous one, but it utilized interaction scenarios without numerical payoffs (e.g., ”If both remain silent, they get minor sentences,” to describe the outcome of mutual cooperation in PD).
+Therefore, the autoformalizing LLM had to infer the appropriate payoffs correctly.
+Since it was not possible to provide a ground truth matrix in this case, we validated the semantic correctness by performing constraint checks (described in Sect.
+3.2) to determine whether the payoff matrix satisfied the constraints for the ground truth game type (Sect.
+2.1).
+4.4 Axelrod’s Tournament
 Strategy Description
@@
 tit-for-tat Start with a default move. Then, mirror the opponent’s move in the previous round.
-Table 5. Strategies utilised in the evaluation.
-In the second experiment, we selected five agents created in Experiment 1, one for each of the five games.
-For each game, we conducted a tournament based on the design of Axelrod’s tournament Axelrod and Hamilton (1981).
-In this setup, six copies of each agent were created and assigned one of six strategies; each agent then played against every other agent, including itself.
+Table 4: Strategies utilized in the evaluation.
+To demonstrate the framework’s usefulness in analysing interaction scenarios, we selected five agents created during the first experiment, one for each of the five games.
+For each game, we conducted a tournament modelled after Axelrod’s tournament Axelrod and Hamilton (1981).
+In this setup, six copies of each agent were created and assigned one of six strategies.
+Each agent then played against every other agent, including itself.
 The strategies used are listed in Table 4.
-To adapt the strategies to games without explicit cooperation and defection concepts, we defined a “default move” and its “opposite” as the basis for certain strategies.
-The default moves were: “Cooperate” for the Prisoner’s Dilemma, “Dove” for Hawk-Dove, “Stag” for Stag Hunt, “Opera” for the Battle of the Sexes, and “Heads” for Matching Pennies.
 The main metric for evaluating strategy performance was the total payoff achieved by an agent using a given strategy in each game.
-To allow for cross-game comparisons of strategy effectiveness, total payoffs were normalized with the following formula:
+To allow for cross-game comparisons of strategy effectiveness, total payoffs were normalized using:
 $norm\_payoff=\frac{total\_payoff-min\_total\_payoff}{{max\_total\_payoff-min\_total\_payoff}},$
 $Total\_payoff$ refers to the sum of the payoffs an agent collects across all rounds and $min\_total\_payoff$ and $max\_total\_payoff$ represent the minimum and maximum total payoffs achievable, respectively.
-4.4. Exp. 3: Autoformalization of Strategies
+4.5 Autoformalization of Strategies
 We then evaluated the semantic correctness of the autoformalized strategies.
-To achieve this, we used the natural language description of the tit-for-tat strategy, along with its Prolog implementation, as a reference example.
+To achieve this, we used the natural language description of the tit-for-tat strategy, along with its Prolog implementation, as an example.
 For the other five strategies, the autoformalization module was provided with their natural language descriptions, similar to those shown in Table 4.
@@
 For the random strategy, correctness was also manually validated.
-This tournament setup was repeated five times for each strategy to ensure consistency in results.
-5. Results and Discussion
-5.1. Exp. 1: Autoformalization of Game Descriptions
-Attempts 1 2 3 4 5
-Count 93% 7% 0 0 12%
-Table 6. Distribution of autoformalization attempts number for 275 agents.
-Table 6 shows the distribution of attempts needed to generate syntactically correct code for the 275 agents created in the first experiment.
-For 93% of agents, syntactically correct code was generated on the first attempt.
-When the correct code was not generated initially, 7% of agents succeeded on the second attempt, while 12% could not produce correct code within the allowed attempts.
-This latter case highlights a need for further refinement of the feedback loop to improve success rates.
-However, the examination of the non-aggregated distribution of the invalid agents showed that an agent without syntactically valid game program was generated in no more than 1 in 5 instances.
-Figure 2. The percentage of syntactic and semantic correctness of generated game rules for each game.
-The overall syntactic correctness rate was 96%, while semantic correctness reached 87%, closely aligning with the results of our previous work (98% and 88%, respectively).
-Figure 2 depicts the distribution of syntactic and semantic correctness across various games.
-The Battle of the Sexes and Stag Hunt emerged as the most error-prone scenarios, exhibiting the lowest semantic correctness rates, though still exceeding 80%.
-The total-payoff-based approximate validation demonstrated 88% semantic correctness, while the exact individual-payoff-based validation achieved 87%.
-Five instances were incorrectly labelled as semantically correct by the approximate method, indicating that the results of the approximate and exact methods were closely aligned.
-A comparison of constraint checking and exact semantic validation revealed a few cases where the autoformalized payoff matrix satisfied the constraints, but errors in other parts of the program – such as role assignment – resulted in incorrect payoff assignments.
-This demonstrates that exact validation is feasible only when the ground truth payoff matrix is available.
-5.2. Exp. 2: Axelrod’s Tournament
-Figure 3. Normalized total payoff for each strategy by game.
-Figure 3 presents the normalized total payoffs for each strategy across all games, while Table 7 provides the average normalized total payoff for each strategy across games.
+This tournament setup was repeated five times.
+5 Results and Discussion
+5.1 Autoformalization of Interaction Descriptions
+1 2 3 4 5
+Numeric payoffs
+Claude 3.5 S 100.00% 0.00% 0.00% 0.00% 0.00%
+GPT-4o 96.77% 2.87% 0.00% 0.00% 0.36%
+Non-numeric payoffs
+Claude 3.5 S 100.00% 0.00% 0.00% 0.00% 0.00%
+GPT-4o 97.64% 2.02% 0.34% 0.00% 0.00%
+Table 5: Distribution of attempts at creating syntactically correct code for each model.
+Table 5 shows the distribution of attempts needed to generate syntactically correct code for both models.
+The type of description (numeric vs.
+non-numeric) had no influence on syntactic correctness of the generated code.
+Claude 3.5 Sonnet achieved syntactically correct code in 100% of cases and GPT-4o in nearly 100%.
+The few errors observed involved non-Prolog comment characters and non-escaped special characters in strings, all of which were successfully corrected on the second attempt.
+This demonstrates that providing a specific line that generated the error in the feedback prompt significantly enhances an LLM’s ability to self-correct compared to merely providing an error message Mensfelt et al.
+(2024).
+The only case where GPT-4o was unable to correct a syntactic error involved the error message “Unexpected end of file“ which indicated the end of the file as the error’s source.
+However, the actual line generating the error was not provided in the feedback prompt.
+Syntactic Runtime Semantic
+Numeric payoffs
+Claude 3.5 Sonnet 100.0% 88.0% 86.0%
+GPT-4o 100.0% 90.0% 86.0%
+Non-numeric payoffs
+Claude 3.5 Sonnet 100.0% 79.0% 67.0%
+GPT-4o 100.0% 91.0% 68.0%
+Table 6: Syntactic, runtime, and semantic correctness for both models and types of scenario descriptions.
+The type of description had no influence on runtime errors in GPT-4o (runtime correctness of 90% and 91% for numeric and non-numeric payoffs, respectively, Table 6).
+However, it did have an influence in the case of Claude 3.5 Sonnet, with a 9 percentage point decrease to 79% for non-numerical payoffs.
+A runtime error in this context means that the program was syntactically correct; however, while querying a predicate during runtime—most often to calculate payoffs—a predicate necessary to execute the query was missing.
+This highlights the need to extend the autoformalization feedback loop to the runtime stage and leverage the tracing mechanism of the solver to identify and create missing predicates.
+Numeric Payoffs
+BS HD MP PD SH
+Syntactic 100.0% 100.0% 100.0% 100.0% 100.0%
+Runtime 89.1% 96.4% 68.5% 96.4% 87.3%
+Semantic 89.1% 96.4% 68.5% 89.1% 87.3%
+Non-Numeric Payoffs
+BS HD MP PD SH
+Syntactic 100.0% 100.0% 100.0% 100.0% 100.0%
+Runtime 82.1% 85.7% 74.0% 70.7% 83.1%
+Semantic 82.1% 77.8% 34.2% 64.0% 83.1%
+Table 7: Correctness by game for Claude 3.5 Sonnet for game descriptions with numeric and non-numeric payoffs.
+Numeric Payoffs
+BS HD MP PD SH
+Syntactic 98.3% 100.0% 100.0% 100.0% 100.0%
+Runtime 83.1% 89.1% 96.4% 89.1% 90.9%
+Semantic 79.7% 89.1% 90.9% 85.5% 83.6%
+Non-Numeric Payoffs
+BS HD MP PD SH
+Syntactic 100.0% 100.0% 100.0% 100.0% 100.0%
+Runtime 93.2% 88.1% 87.3% 96.5% 91.5%
+Semantic 93.2% 67.8% 30.2% 73.7% 76.3%
+Table 8: Correctness by game for GPT-4o for game descriptions with numeric and non-numeric payoffs.
+The semantic correctness was lower than runtime correctness for both models and description types (Table 6).
+This means that even though an agent was able to select a move during runtime and update its state, the semantics of the program were not correct.
+For descriptions with numerical payoffs, this implies that the payoff matrix does not contain the required payoffs.
+For descriptions without numerical payoffs, it indicates that the payoff matrix does not satisfy the constraints of the required game.
+Tables 7 and 8 present the syntactic, runtime, and semantic correctness for different games.
+The runtime and semantic correctness varied depending on the model and the game.
+Notably, Matching Pennies consistently resulted in low semantic correctness for non-numerical descriptions, with 34.2% correctness for Claude 3.5 Sonnet and 30.2% correctness for GPT-4o.
+This may be attributed to some descriptions specifying which player wins if the choices match or mismatch, without explicitly stating that the other player will lose—a constraint of the game.
+Extending the autoformalization feedback loop to include semantic evaluation has the potential to significantly improve correctness.
+5.2 Axelrod’s Tournament
+Figure 2: Normalized total payoff for each strategy by game.
+Figure 2 presents the normalized total payoffs for each strategy across all games.
 Overall, the best-response strategy proved to be the most effective on average, followed closely by tit-for-tat.
@@
 In the Prisoner’s Dilemma, best-response achieved a lower payoff than the anti-default-move strategy (analogous to “defect”), as it initiated with the “cooperate” move.
-The results demonstrate that agents with autoformalized rules can effectively simulate and assess the performance of various strategies within different game-theoretic scenarios.
-Strategy Payoff [%]
-anti-default-move 50.90
-anti-tit-for-tat 55.22
-best-response 84.08
-default-move 73.96
-random 61.08
-tit-for-tat 79.84
-Table 7. Average normalized total payoffs for each strategy.
-5.3. Exp. 3: Autoformalization of Strategies
-Strategy Correctness [%]
-anti-default-move 100.00
-anti-tit-for-tat 100.00
-best-response 40.00
-default-move 100.00
-random 60.00
-Table 8. Percentage of correctness for each strategy.
+The results demonstrate that agents with autoformalized rules can effectively simulate and assess the performance of various strategies within different modelled scenarios.
+5.3 Autoformalization of Strategies
+Strategy Claude 3.5 Sonnet GPT-4o
+anti-default-move 100.00% 100.00%
+anti-tit-for-tat 100.00% 100.00%
+best-response 80.00% 100.00%
+default-move 100.00% 100.00%
+random 80.00% 100.00%
+Table 9: Percentage of semantic correctness for each strategy.
 Table 9 presents the percentage of correctly autoformalized strategies.
-The simpler strategies—default-move, anti-default-move, and anti-tit-for-tat—were always correctly autoformalized.
-In contrast, the random strategy achieved correct autoformalization in 60% of cases, and best-response, which requires selecting the move that maximizes payoff in response to the opponent’s previous move, was correctly autoformalized in 40% of cases.
-Since all strategies were generated using only the tit-for-tat implementation as an example, increasing the diversity of examples provided to the autoformalization module could potentially improve its accuracy in correctly formalising more complex strategies.
-6. Related Work
+GPT-4o achieved 100% correctness in autoformalizing all five strategies.
+Claude 3.5 Sonnet produced one incorrect strategy for best-response and one for random, which are the strategies that deviate the most from the tit-for-tat strategy presented as an example in the prompt.
+One error occurred during runtime, and another during semantic validation.
+Overall, the high correctness demonstrates that the GAMA framework has the potential to effectively autoformalize both game rules and gameplay strategies.
+6 Related Work
 Originally proposed for agents in game scenarios, the Game Description Language (GDL) offers a promising approach for formalising complex strategic interactions in multi-agent environments by providing a declarative framework to define game rules, legal moves, and outcome conditions Genesereth et al.
@@
 For instance, Yang et al.
-Yang et al.
 (2023) applied GPT-3 with few-shot prompting to translate natural language sentences into formal representations, which were then inputted into answer set programs.
 While effective on NLP benchmarks, this method faced occasional translation errors.
-To further enhance performance, Pan et al.Pan et al.
+To further enhance performance, Pan et al.
 (2023) combined in-context learning with self-refinement based on solver evaluations, surpassing standard LLM and chain-of-thought prompting approaches on logical reasoning benchmarks.
-In the context of temporal logic, Cosler et al.Cosler et al.
+In the context of temporal logic, Cosler et al.
 (2023) addressed the problem of translation errors with an interactive system allowing users to refine sub-translations.
@@
 (2023).
-7. Conclusions and Future Work
-We have presented a framework for the autoformalization of game-theoretic scenarios, specifically those modelled by $2\times 2$ simultaneous-move games.
-The framework also enables autoformalization of strategies for these games and the simulation of tournaments based on the autoformalized rules.
-The generated programs undergo three levels of validation: syntactic validation by consulting a solver, functional testing through gameplay, and semantic validation – either approximate or exact – by comparing achieved payoffs with target outcomes.
-Evaluation of the framework demonstrated high syntactic and semantic accuracy of the code generated by agents across various interaction scenarios.
+7 Conclusions and Future Work
+We have presented a framework for the autoformalization of interaction scenarios, specifically those modelled by $2\times 2$ simultaneous-move games.
+The framework enables the autoformalization of strategies for these games and the simulation of tournaments based on the autoformalized rules.
+The generated programs undergo three levels of validation: syntactic validation by consulting a solver, functional testing through gameplay, and semantic validation by comparing achieved payoffs with target outcomes or by checking the game constraints.
+Evaluation of the framework demonstrated high syntactic and semantic accuracy in the code generated by agents across various interaction scenarios.
 Future work will focus on extending the framework to model games beyond $2\times 2$ simultaneous-move scenarios.
 Our previous work Mensfelt et al.
-(2024) demonstrated that our sample solver for the Prisoner’s Dilemma could successfully serve to generate code for structurally similar games, such as rock-paper-scissors (with a different number of moves) and sequential versions of the Prisoner’s Dilemma.
+(2024) demonstrated that using our sample solver for the Prisoner’s Dilemma we could successfully generate code for structurally similar games, such as rock-paper-scissors (with a different number of moves) and sequential versions of the PD.
 However, achieving successful generalization to a broader class of games will require assembling a more diverse set of in-context learning examples and developing methods to retrieve the most relevant examples for autoformalization in specific scenarios.
-The flexibility and generalisability of the introduced framework also make it applicable to games and game-like interactions beyond game theory.
-Another direction for future work involves enhancing the feedback loop to more effectively leverage a large language model.
-At present, the LLM receives feedback solely from the trace generated when consulting the solver with the produced code.
-In the next phase, we plan to expand this feedback mechanism by incorporating runtime errors, thereby creating a more robust autoformalization module.
-Finally, in cases where unsupervised autoformalization proves infeasible for complex scenarios, we plan to introduce an interactive mode.
-This mode will allow users to iteratively refine the generated code alongside the autoformalization module.
-These interactively generated programs will then be added to our existing program library, expanding the resources available for in-context learning and enabling the autoformalization of increasingly complex scenarios.
-{acks}
+Another direction for future work involves extending the autoformalization feedback loop to include runtime and semantic validation checks, further improving the correctness of the generated programs.
+These extensions will provide a robust framework for generating reasoning modules for decision-making agents in a wider class of games and game-like interactions.
+Acknowledgments
 This work was supported by a Leverhulme Trust International Professorship Grant (LIP-2022-001).

### v3

@@
-GAMA: Generative Agents for Multi-Agent Autoformalization
+Generative Agents for Multi-Agent Autoformalization of Interaction Scenarios
 Abstract
-Multi-agent simulations facilitate the exploration of interactions among both natural and artificial agents.
-However, modelling real-world scenarios and developing simulations often requires substantial expertise and effort.
-To streamline this process, we present a framework that enables the autoformalization of interaction scenarios using agents augmented by large language models (LLMs) utilising game-theoretic formalisms.
-The agents translate natural language descriptions of interactions into executable logic programs that define the rules of each game, ensuring syntactic correctness through validation by a solver.
-A tournament simulation then tests the functionality of the generated game rules and strategies.
-After the tournament, if a ground truth payoff matrix is available, an exact semantic validation is performed.
-We evaluate our approach on a diverse set of 110 natural language descriptions exemplifying five $2\times 2$ simultaneous-move games, achieving 100% syntactic and 76.5% semantic correctness in the generated game rules for Claude 3.5 Sonnet, and 99.82% syntactic and 77% semantic correctness for GPT-4o.
-Additionally, we demonstrate high semantic correctness in autoformalizing gameplay strategies.
-Overall, the results highlight the potential of autoformalization to leverage LLMs in generating formal reasoning modules for decision-making agents.
+Multi-agent simulations are versatile tools for exploring interactions among natural and artificial agents, but their development typically demands domain expertise and manual effort.
+This work introduces the Generative Agents for Multi-Agent Autoformalization (GAMA) framework, which automates the formalization of interaction scenarios in simulations using agents augmented with large language models (LLMs).
+To demonstrate the application of GAMA, we use natural language descriptions of game-theoretic scenarios representing social interactions, and we autoformalize them into executable logic programs defining game rules, with syntactic correctness enforced through a solver-based validation.
+To ensure runtime validity, an iterative, tournament-based procedure tests the generated rules and strategies, followed by exact semantic validation when ground truth outcomes are available.
+In experiments with 110 natural language descriptions across five $2\times 2$ simultaneous-move games, GAMA achieves 100% syntactic and 76.5% semantic correctness with Claude 3.5 Sonnet, and 99.82% syntactic and 77% semantic correctness with GPT-4o.
+The framework also shows high semantic accuracy in autoformalizing agents’ strategies.
 1 Introduction
-Multi-agent simulations provide insights into diverse phenomena, ranging from individual human interactions and organizational dynamics to biological processes, robotics, and interactions among artificial agents.
+Multi-agent simulations (MAS) provide insights into diverse phenomena, ranging from individual human interactions and organizational dynamics to biological processes, robotics, and interactions among artificial agents.
 However, modelling real-world scenarios and implementing corresponding simulations typically require substantial human expertise and effort.
-Consider, for instance, the challenge of coordinating autonomous vehicles: creating a simulation model involves specifying the rules of interaction, the goals of each agent, the information available to them, and the strategies they may employ, among other factors.
-Once developed, this model must then be implemented.
-The emergence of Large Language Models (LLMs) enables a novel method for streamlining these tasks through autoformalization.
-We demonstrate it using scenarios that can be modelled as $2\times 2$ simultaneous-move games.
-Although LLMs have been evaluated as agents in game-theoretic contexts Akata et al.
-(2023); Fan et al.
-(2023); Guo (2023); Lorè and Heydari (2023), their inherent limitations—such as hallucinations Achiam et al.
-(2023) and logical or arithmetical errors Imani et al.
-(2023)—constrain their effectiveness as rational decision—makers in simulations.
-Rather than relying on LLMs as autonomous agents, a more promising approach is to harness their strengths in format translation, particularly in converting natural language into formal representations.
-This process, known as autoformalization Wu et al.
-(2022), has shown promise in formalizing mathematics Wu et al.
-(2022); He-Yueya et al.
-(2023); Jiang et al.
-(2022) and logic Yang et al.
-(2023); Pan et al.
-(2023); Cosler et al.
-(2023); Chen et al.
-(2023); Feng et al.
-(2023).
-The resulting formal representations can be utilized in combination with subsymbolic models within a neuro-symbolic framework.
-This approach, as demonstrated by recent advances such as AlphaGeometry, has the potential to achieve human-like problem-solving abilities in challenging domains like geometry Trinh et al.
-(2024).
-In this work, we introduce GAMA: Generative Agents for Multi-Agent Autoformalization, a framework that leverages autoformalization to generate formal representations of strategic interactions.
-These representations serve as the reasoning module of agents deployed in simulations.
-This research builds upon our prior work Mensfelt et al.
-(2024), where we presented a basic framework for autoformalizing natural language interaction scenarios.
-In this paper, we integrate the previously developed autoformalization module into a modular agent architecture within a simulation framework.
+Recent advances in large language models (LLMs) and autoformalization techniques [37] offer a promising new approach to streamlining these tasks.
+In this work, we demonstrate this approach using interaction scenarios modelled as $2\times 2$ simultaneous-move games.
+LLMs have been increasingly explored as agents in interaction scenarios and game-theoretic contexts, though often under a weak notion of agency [2, 7, 14, 19].
+Furthermore, their inherent limitations—such as hallucinations [1] and logical or arithmetic errors [16]—undermine their reliability as rational decision-makers in simulations.
+Addressing this challenge, autoformalization shifts the role of LLMs from decision-making to format translation, converting natural language interaction descriptions into formal representations.
+This leverages one of LLMs’ core strengths, with autoformalization showing considerable promise in formalizing mathematics [37, 15, 17] and logic [38, 26, 6, 5, 8].
+In multi-agent systems, such formalized interaction descriptions can aid and, often act, as decision-making modules, reducing reliance on the LLM’s reasoning correctness during runtime.
+Moreover, this formal representation can be iteratively refined using automatic reflection and refinement mechanisms [21, 29, 32, 13], enabling continuous improvement via feedback loops.
+This shift from natural language reasoning to formal, verifiable representations establishes a more robust foundation for automating LLM-driven agent design.
+Our prior work [23], introduced an autoformalization module capable of converting interaction descriptions into syntactically validated code.
+The Generative Agents for Multi-Agent Autoformalization (GAMA) incorporates the previous work as a component of the simulation framework, while additionally providing runtime and semantic validation and enabling the generation of formal representations of both strategic interactions and game strategies.
+These formalizations act as reasoning modules for agents deployed in simulations, enabling comparison of different strategies for given interactions.
 The main contributions of this work are as follows:
   1. I.
-    An agent model with formal game and strategy representation: We propose an agent model that incorporates formal representations of both games and strategies, along with the autoformalization module from our earlier work.
-Game rules and strategies can either be autoformalized or provided using predefined rules.
+    Agent model with formal game and strategy representations: We propose an agent architecture that integrates formalized representations of games and strategies alongside the autoformalization module.
+Game rules and strategies can be autoformalized or supplied as predefined specifications.
   2. II.
-    Three-level agent validation: We introduce a validation mechanism for agents at three levels: (1) syntactic correctness, using formal solvers; (2) functionality, by testing the code in the simulation; and (3) semantic correctness, by comparing target and achieved payoffs (when a ground truth payoff matrix is available) or verifying compliance with game type constraints (when a ground truth game type is provided).
+    Three-level agent validation: We introduce a multi-level validation pipeline for agents, including (1) syntactic validation using formal solvers, (2) runtime validation through simulation-based testing, and (3) semantic validation based on ground truth outcomes (when available) or compliance with game-specific constraints.
   3. III.
-    A simulation framework for autoformalized scenarios: We present a simulation framework for autoformalized interaction scenarios, enabling automated validation and facilitating the exploration and comparison of different strategies for specific scenarios.
+    Simulation framework for autoformalized scenarios: We present a simulation environment that supports automated validation and facilitates the exploration and comparison of strategies formalized from natural language descriptions.
   4. IV.
-    Experimental evaluation: We evaluate the framework, demonstrating high levels of syntactic and semantic correctness for scenarios modelled as $2\times 2$ simultaneous-move games.
-Additionally, we assess the agents’ ability to autoformalize five distinct strategies.
+    Experimental evaluation: We empirically evaluate GAMA, demonstrating high syntactic and semantic correctness across 110 scenarios modelled as $2\times 2$ simultaneous-move games.
+Furthermore, we assess the agents’ ability to autoformalize and execute five distinct gameplay strategies.
+The remainder of this paper is organized as follows.
+Section 2 introduces the background and preliminaries.
+Section 3 presents the GAMA framework.
+Section 4 details the experimental setup and evaluation methodology.
+Section 5 reports the experimental results.
+Finally, Section 6 concludes the paper and outlines directions for future work.
 2 Preliminaries
 2.1 Game Theory
-Game theory is an analytical framework for strategic interactions, where games are composed of four main components Rasmusen (2006):
-  • Players: players consistently and willingly act towards maximising their utility, in full awareness of the situation they are involved in;
-  • Actions: mappings between states and outcomes, carried out players Osborne and Rubinstein (1994);
-  • Payoffs: payoffs are quantitative measures over the consequent outcomes of action tuples chosen by players Gibbons (1992);
-  • Information: game information describes a player’s knowledge about the available actions, historical moves, and the player’s image of its opponents Osborne and Rubinstein (1994).
+Game theory is an analytical framework for strategic interactions, where games are composed of four main components [28]:
+Players
+players consistently and willingly act towards maximising their utility, in full awareness of the situation they are involved in.
+Actions
+mappings between states and outcomes, carried out by players [25].
+Payoffs
+payoffs are quantitative measures over the consequent outcomes of action tuples chosen by players [11].
+Information
+game information describes a player’s knowledge about the available actions, historical moves, and the player’s image of its opponents [25].
 We formally define games in this context as a tuple $(N,A,\Pi)$ consisting of a set of $n$ players $N$, a non-empty set of actions $A_{i},\forall i\in N$, and for each player $i$ a payoff function $\pi_{i}:\times_{j\in N}A_{j}\rightarrow\mathbb{R}$ mapping action profiles to player payoffs.
+Table 1: A general normal form game describing the four possible outcomes $W,X,Y,Z$, where the tuples denote the row player’s and column player’s corresponding payoffs, in general: $(\pi_{row},\pi_{col})$.
 row/col L R
@@
 D ($Y_{row}$, $Y_{col}$) ($Z_{row}$, $Z_{col}$)
-Table 1: A general normal form game describing the four possible outcomes $W,X,Y,Z$, where the tuples denote the row player’s and column player’s corresponding payoffs, in general: $(\pi_{row},\pi_{col})$.
 In the case of symmetric games, the payoffs solely depend on the chosen strategies, while with asymmetric games, the outcomes yield different payoffs based on the player’s identity.
 In our experiments, the Prisoner’s Dilemma, Hawk-Dove, and Stag-Hunt are symmetric, and we classify actions U and L as “Cooperate” denoted by $C$ and actions D and R as “Defect” denoted by $D$.
-This allows us to use the terminology from Axelrod’s tournaments Axelrod and Hamilton (1981) to define these games in terms of the four outcomes: $\mathbf{T}:(D,C),\mathbf{R}:(C,C),\mathbf{P}:(D,D),\mathbf{S}:(C,D)$.
-  • Prisoner’s Dilemma is characterised by a dominant temptation to defect, represented by the payoff relation $T>R>P>S$,
-  • Hawk-Dove entails switched relative payoffs between the punishment for mutual defection and the sucker’s payoff – $T>R>S>P$,
-  • Stag Hunt involves a relatively high reward for mutual cooperation over the temptation to defect – $R>T>P>S$.
+This allows us to use the terminology from Axelrod’s tournaments [4] to define these games in terms of the four outcomes: $\mathbf{T}:(D,C),\mathbf{R}:(C,C),\mathbf{P}:(D,D),\mathbf{S}:(C,D)$.
+Prisoner’s Dilemma
+is characterised by a dominant temptation to defect, represented by the payoff relation $T>R>P>S$.
+Hawk-Dove
+entails switched relative payoffs between the punishment for mutual defection and the sucker’s payoff: $T>R>S>P$.
+Stag Hunt
+involves a relatively high reward for mutual cooperation over the temptation to defect: $R>T>P>S$.
 The Battle of the Sexes and Matching Pennies are asymmetric games.
 As the outcomes mean different payoffs for row and column, we cannot use the generalization from above and have to define the games from both players’ perspectives:
-  • Battle of the Sexes is a coordination game, where both players prefer an agreement but have different preferences over the coordinated outcomes:
-    • row: $W>Z>\{X,Y\}$,
+Battle of the Sexes
+is a coordination game, where both players prefer an agreement but have different preferences over the coordinated outcomes:
+  • Row: $W>Z>\{X,Y\}$.
   • Column: $Z>W>\{X,Y\}$.
-  • Matching Pennies: coordinated outcomes reward the row player, while mismatched actions benefit the column player:
-    • row: $\{W,Z\}>\{X,Y\}$,
+Matching Pennies
+coordinated outcomes reward the row player, while mismatched actions benefit the column player:
+  • Row: $\{W,Z\}>\{X,Y\}$.
   • Column: $\{X,Y\}>\{W,Z\}$.
 2.2 Large Language Models
-The rapid advancement of natural language processing (NLP), driven by transformer architectures Gillioz et al.
-(2020) and pre-trained models Qiu et al.
-(2020), has led to the emergence of Large Language Models (LLMs).
-State-of-the-art (SOTA) LLMs continue to evolve rapidly, primarily due to the increasing scale of model parameters and training datasets Zhao et al.
-(2023).
-In this work, we employ OpenAI’s GPT-4 Omni OpenAI (2024), a multimodal SOTA model capable of processing text, images, and audio.
-It has demonstrated strong performance on reasoning benchmarks Wang and Zhao (2024).
-Likewise, we utilize Anthropic’s Claude 3.5 Sonnet, another high-performing SOTA model, which excels across multiple benchmarks Anthropic (2024).
+The rapid advancement of natural language processing (NLP), driven by transformer architectures [12] and pre-trained models [27], has led to the emergence of Large Language Models.
+State-of-the-art (SOTA) LLMs continue to evolve rapidly, primarily due to the increasing scale of model parameters and training datasets [39].
+In this work, we employ OpenAI’s GPT-4 Omni [24], a multimodal SOTA model capable of processing text, images, and audio.
+It has demonstrated strong performance on reasoning benchmarks [35].
+Likewise, we utilize Anthropic’s Claude 3.5 Sonnet, another high-performing SOTA model, which excels across multiple benchmarks [3].
 2.3 General Game Playing
-General game playing Genesereth et al.
-(2005) focuses on developing intelligent systems that explicitly represent the rules of arbitrary new games and learn to play them autonomously, without human intervention.
-The Game Description Language (GDL) has been proposed as a formal, machine-processable language for describing the rules of arbitrary games Love et al.
-(2006).
-GDL focused on information games only, so it was extended in GDL-II Thielscher (2010) to cover n-player games with incomplete information and games in extensive normal form Thielscher (2011).
+General game playing [9] focuses on developing intelligent systems that explicitly represent the rules of arbitrary new games and learn to play them autonomously, without human intervention.
+The Game Description Language (GDL) has been proposed as a formal, machine-processable language for describing the rules of arbitrary games [20].
+GDL focused on information games only, so it was extended in GDL-II [33] to cover n-player games with incomplete information and games in extensive normal form [34].
 A challenge with GDL systems is that learning without human guidance demands complex reasoning.
 Players must infer the possible actions of others, effectively analysing hypothetical game scenarios before making decisions.
-Action formalisms like the classical Situation Calculus McCarthy and Hayes (1981) have been developed for precisely this purpose.
-Formal schemes and inference methods are readily available for Situation Calculus Giacomo et al.
-(2010); Lesperance et al.
-(2024), while their deployment in general game playing presupposes a translation from GDL into existing, suitably expressive action languages.
-One such scheme Schiffel and Thielscher (2011) shows how to fully embed GDL-II into a version of the Situation Calculus based on knowledge fluents Scherl and Levesque (2003).
+Action formalisms like the classical Situation Calculus [22] have been developed for precisely this purpose.
+Formal schemes and inference methods are readily available for Situation Calculus [10, 18], while their deployment in general game playing presupposes a translation from GDL into existing, suitably expressive action languages.
+One such scheme [31] shows how to fully embed GDL-II into a version of the Situation Calculus based on knowledge fluents [30].
 Our game solver is inspired by these previous works, to support light GDL specifications that act as target representation for the autoformalization.
@@
 3.1 Overview
-Figure 1: GAMA framework: each agent’s game rules and strategy can be predefined, autoformalized, or loaded from an existing agent.
-The initial agent pool undergoes syntactic validation, and agents that fail to produce syntactically valid code within a defined number of attempts are eliminated (denoted by a crossed circle).
-The remaining agents are validated in the tournament and subsequently undergo semantic validation.
-The GAMA framework introduced in this work enables the creation of modular autoformalizing agents and supports tournament play using these agents (Fig.
-1).
-An agent consists of an autoformalization module and a logic programming component implemented in Prolog, which is divided into two modules: a Game module, containing both game-independent and game-dependent rules, and a Strategy module, which describes the actions to select against an opponent.
+Figure 1: Overview of the GAMA framework.
+Dashed lines indicate optional control flows.
+Games and strategies can be specified either through predefined rules or natural language descriptions, or optionally by loading a predefined agent.
+Semantic validation is performed when ground truth outcomes are available.
+The Validator may employ a formal solver.
+Fig.1 presents an overview of the GAMA framework.
+The framework is implemented in Python and Prolog [36], though its modular architecture supports substituting alternative solvers.
+The source code and log files generated during the evaluation are available at [^1] https://github.com/dicelab-rhul/GAMA.
+Each agent comprises an autoformalization module and a logic programming component, which is further divided into two modules: a Game module, containing both game-independent and game-specific rules, and a Strategy module, which specifies the agent’s action selection logic against an opponent.
 Together, these modules form a solver that encapsulates the game logic and determines the agent’s next move.
-A Python wrapper facilitates the agent’s interaction with the external environment, manages the game history (tracking each agent’s moves, opponent moves, and payoffs), and updates the solver’s state.
-This modular design allows for easy substitution of game-dependent rules and strategies, either through predefined configurations or by autoformalizing natural language descriptions.
-The source code and log files generated during the evaluation are available at [^1] https://github.com/dicelab-rhul/GAMA.
-3.2 Solver
-The solver comprises a game-independent part that represents the rules for any extensive-form game, a game-dependent part that defines the rules of a particular game using the predicates from the game-independent part, and a strategy part to select an action.
-Our solver is implemented in Prolog Körner et al.
-(2022); Wielemaker et al.
-(2012)—a declarative programming language well-suited for high-level problem definitions—but it can easily be substituted with another formal solver.
-The game-independent and game-dependent parts were described in detail in Mensfelt et al.
-(2024).
-To enable the agent incorporating the solver to play, we extended it with the strategy part.
-Additionally, we implemented semantic validation in Prolog using validity constraints.
-The strategy part and constraint-checking mechanism are described below.
-3.2.1 Strategy
-Strategies are submitted separately from the game-dependent component.
-The select/4 predicate is used to determine the next move M of player P against opponent O in game situation S, as illustrated in the tit-for-tat strategy example below:
-select(P, O, S, M) if
-not holds(last_move(O, _LMo), S) and
-holds(default_move(P, M), S).
-select(P, O, S, Mo) if
-holds(last_move(O, Mo), S).
-Here, a player P selects the default move first (typically to cooperate), otherwise it mirrors the opponent’s move in the last round.
-holds/2 is based on the implementation of the Situation Calculus McCarthy and Hayes (1981).
-The strategy relies on specific predicates – default_move/2 and last_move/2 – which are required from the game-dependent part of the solver.
-More sophisticated strategies can be expressed.
-3.2.2 Validity checking
-We check that the auto-formalized payoff matrix satisfies the structural and content constraints of the specified game (as outlined in Sect.
-2.1), by defining validity constraints.
-The listing below specifies the payoff matrix validity constraints for the Prisoner’s Dilemma.
-valid_pd_payoffs(T,R,P,S,C,D) if
-payoff(C,C,R,R) and
-payoff(C,D,S,T) and
-T>R and
-payoff(D,C,T,S) and
-payoff(D,D,P,P) and
-R>P and
-P>S.
-Since the autoformalization module does not impose constraints on the names of actions, this predicate also identifies which actions names correspond to specific actions.
-3.3 Autoformalization Module
+A Python wrapper enables interaction with the external environment, manages game history (including each agent’s moves, opponent actions, and payoffs), and updates the solver’s internal state accordingly.
+3.2 Agents’ Initialization and Syntactic Validation
 Algorithm 1 Generating game rules predicates from natural language descriptions. PD stands for Prisoner’s Dilemma.
@@
 17:return Unable to generate valid predicates within maximum attempts.
-To perform autoformalization, a large language model is employed.
-The framework supports the use of any language model by providing an abstract interface.
-Autoformalization of a natural language description of a game or strategy is achieved through one-shot prompting, using game-dependent predicates for the Prisoner’s Dilemma or the tit-for-tat strategy as examples, respectively.
-After generating the code, its syntactic correctness is validated.
-If errors are detected, the LLM receives feedback, including the erroneous lines, error messages, and instructions for correcting the predicates.
-The max_attempts parameter specifies the maximum number of correction attempts.
-The overview of autoformalization algorithm is presented in Listing 1 Mensfelt et al.
-(2024).
-3.4 Simulation
-The framework is employed in this work in two primary modes: validation and simulation.
-These modes facilitate the evaluation of both the semantic correctness of generated programs and the effectiveness of strategies within various game scenarios.
-In validation mode, the framework ensures the semantic correctness of generated programs by comparing their performance against predefined criteria.
-Once a scenario is autoformalized, an agent is created and provided with a specified strategy.
-The agent then plays against a clone of itself for a set number of rounds.
-If a ground truth payoff matrix is available, the framework performs semantic validation by verifying that the payoff for each unique combination of moves matches the target payoff for that combination.
-Additionally, if the ground truth type of the game is known, the framework enforces constraint checking to confirm that the payoff matrix adheres to all specified constraints for the given game.
-In simulation mode, the framework evaluates and compares strategies to determine their effectiveness within a game scenario.
-Agents equipped with different strategies compete in a round-robin tournament, where each agent plays against every other.
-By simulating multiple rounds, the framework identifies the strategy that performs best on average, offering insights into the relative strengths and weaknesses of competing approaches.
-The parameters of the tournament are as follows:
-  • Agent Pool: Contains all agents participating in the tournament. Each agent can be equipped with game rules and strategies either through autoformalization or predefined configurations.
-  • Number of Rounds: Specifies how many rounds each agent will play.
-  • Match Maker: A function that determines the pairing of agents for matches.
-  • Target Payoffs: Optional parameter used for semantic validation.
-The framework provides two methods for determining the tournament’s outcome:
+To perform autoformalization, an agent employs a large language model (LLM) through a unified abstract interface, allowing compatibility with any LLM.
+Natural language descriptions of games or strategies are formalized using one-shot prompting, with examples of game-dependent predicates (e.g., for the Prisoner’s Dilemma or tit-for-tat strategy) provided in the prompt.
+Once the code is generated, its syntactic correctness is validated.
+If errors are detected, the LLM receives feedback containing the erroneous lines, error messages, and instructions for correcting the predicates.
+The process is repeated up to a limit defined by the max_attempts parameter.
+An overview of the autoformalization module [23] is shown in Listing 1.
+3.3 Simulation and Runtime Validation
+After generating a pool of syntactically correct agents, the framework operates in two primary modes: validation and strategy evaluation.
+These modes enable assessment of both the runtime correctness of generated programs and the effectiveness of strategies across various game scenarios.
+The key parameters for tournament-based simulation are as follows:
+Agent Pool
+All participating agents, initialized with game rules and strategies obtained either through autoformalization or predefined configurations.
+Number of Rounds
+The number of rounds each agent pair will play.
+Match Maker
+A customizable function that defines how agents are paired for matches.
+Target Payoffs
+(Optional) Used for semantic validation when ground truth payoffs are available.
+Tournament outcomes can be determined in two ways:
   1. 1.
-    Selecting agents whose total payoff matches their target payoff, applicable when target payoffs are specified (validation).
+    Selecting agents that achieve the highest total payoff, reflecting strategy effectiveness.
   2. 2.
-    Selecting agents that achieve the highest total payoff, suitable for assessing the success of strategies (simulation).
-Furthermore, the customizable Match Maker function allows for experiments with tailored matching protocols
+    Selecting agents whose total payoff matches the target payoff (if available), used primarily for validation.
+In strategy evaluation mode, the framework compares strategies based on their performance within game scenarios.
+Agents equipped with different strategies typically compete in a round-robin tournament, where each agent plays against every other agent for a specified number of rounds.
+However, alternative competition formats can be defined using a custom Match Maker function.
+By simulating multiple matches, the framework identifies which strategies perform best on average, providing insights into their relative strengths and weaknesses.
+3.4 Semantic Validation
+Semantic validation is performed when ground truth outcomes are available.
+Here, “semantic correctness” refers to the correspondence between the generated code and the input natural language description.
+For example, formalized code for the Prisoner’s Dilemma that is syntactically valid and executes correctly in simulation, but encodes an incorrect payoff matrix, would be considered semantically incorrect.
+In semantic validation mode, when specific target outcomes are provided, the framework performs exact validation.
+Once a scenario is autoformalized and an agent is instantiated with the given strategy, the agent plays against a clone of itself for a fixed number of rounds.
+The resulting outcomes for all action tuples are then compared directly to the target outcomes to assess semantic correctness.
+If specific target outcomes are not available but the game type is known (e.g., it is known that the scenario represents a Prisoner’s Dilemma, but the exact payoff matrix is unknown), the framework performs constraint-based validation.
+In this case, it verifies that the generated payoff matrix satisfies all necessary properties (such as symmetry or payoff dominance relations) defined for that game type (Section 3.5.4).
+This ensures that the generated logic faithfully captures the intended structure of the game.
+3.5 Solver
+Our solver is based on logic programming, as described in [23], but repeated here for completeness.
+It comprises of a game-independent part representing the rules for any normal-form game, a game-dependent part defining the rules of a particular game using the predicates from the game-independent part, and a set of auxiliary predicates completing the definition of that particular game.
+We follow standard Prolog conventions to interpret a game: variables are indicated by uppercase letters, while predicates and function symbols by lowercase.
+An underscore ’_’ represents a variable whose value is ignored within a definition.
+The game state is then a situation, initially denoted by a constant (e.g.
+s0).
+The binary function do(M, S) denotes the situation resulting from the executing move M in situation S, consistent with the Situation Calculus.
+3.5.1 Game-independent part
+We define all legal transitions of an extended-form game from an initial situation S to a final situation F as follows:
+game(F, F) if
+final(F).
+game(S, F) if
+not final(S) and
+legal(M, S) and
+game(do(M, S), F).
+The game terminates when the final game situation F is reached.
+Otherwise, in a non-final situation S, the game accepts a legal move M, and the game continues in the next do(M,S) situation, until the final situation F is reached.
+To determine what holds in each legal situation, we use Situation Calculus, represented as:
+holds(F, S) if
+initially(F, S).
+holds(F, do(M, S)) if
+effect(F, M, S).
+holds(F, do(M, S)) if
+holds(F, S) and
+not abnormal(F, M, S).
+holds/2 here is similar to true/1 in GDL, but with an additional parameter of the situation in which the fluent is true.
+It states that a fluent F holds in the initial situation (init/1 in GDL), a new fluent F is initiated by the effects of a move M executed in a situation S (next/1 in GDL), and a fluent F persists after a move is made, provided it is not abnormal; abnormal fluents do not persist (implicit in GDL).
+We also use rules of the form:
+finally(F, S) if Conditions.
+to return derived fluents F describing the result of the game, when the Conditions hold in the final situation S.
+3.5.2 Game-dependent part
+To represent a specific game we need to define game-dependent predicates for the initial state initial/1, the legal moves legal/2, what holds in the initial game situation via initially/2, the effects of a move on a situation via effect/3, what stops persisting in a situation after the execution of a move via abnormal/3, the final situation final/1, and the result of the game via finally/2 definitions.
+To exemplify these definitions, we show how to describe a PD game to our solver.
+The initial situation s0 is defined as:
+initial(s0).
+What holds in this initial situation we specify it as:
+initially(player(p1), s0).
+initially(player(p2), s0).
+initially(role(p1,row), s0).
+initially(role(p2,col), s0).
+initially(control(p1), s0).
+initially(control(p2), s0).
+These assertions define first the player names represented by unique identifiers (p1 and p2), their roles (p1 is the row player, while p2 is the column player), and then the fact that initially either of them can the game next (via the control/1 fluent,as in GDL).
+What holds in the initial situation changes as a result of move being made in the game.
+We represent moves as terms of the form move(P, M), where P is a player, and M is a move.
+As in PD it is possible for any player to defect (’D’) or cooperate (’C’), we express this as:
+possible(move(P,’D’),S) if holds(player(P),S).
+possible(move(P,’C’),S) if holds(player(P),S).
+It is then legal for a player to make a possible move if they have the control to execute it in the current situation:
+legal(move(P, M), S) if
+possible(move(P, M), S) and
+holds(control(P), S).
+As an effect of a legal move M being made by a player P, we record in the next situation that the player did that move (effect/3 is like next/1 in GDL):
+effect(did(P, M), move(P, M), S).
+Once a legal move is executed, the player loses control, which we specify in our framework as:
+abnormal(control(P), move(P, M), S).
+When a player loses control cannot play a move until control is passed back. When each player makes a move from the initial situation, the resulting final situation can be determined as:
+final(do(M2, do(M1, S))) if
+ground(M2) and
+ground(M1) and
+initial(S).
+Assuming the payoff matrix following the structure presented in Table 1 defined as:
+payoff(’D’, D’, 1, 1).
+payoff(’C’, D’, 0, 5).
+payoff(’D’, C’, 5, 0).
+payoff(’C’, C’, 3, 3).
+the outcome of the game holds information about the actual moves made by the players, and their payoffs.
+finally(outcome(P1, M1, U1, P2, M2, U2), S) if
+holds(role(P1, row), S) and
+holds(did(P1, M1), S) and
+holds(role(P2, col), S) and
+holds(did(P2, M2), S) and
+payoff(M1, M2, U1, U2).
+We can then extract specific outcome information, e.g. the player’s utility, through a goal/2 fluent (as in GDL) e.g.:
+finally(goal(P1, U1), S) if
+finally(outcome(P1, _, U1, _, _, _), S).
+finally(goal(P2, U2), S) if
+finally(outcome(_, _, _, P2, _, U2), S).
+This finalises the definition of PD, enabling us to use the game description above to reason about the game’s dynamics.
+For instance, if player p1 wanted to analyse the game and identify the actions required to achieve a utility of 5, then it asks the query:
+?- game(s0,F), finally(goal(p1,5),F).
+Applying the PD payoff matrix, our solver produces two solutions for F, separated by semicolons (;). The result false indicates "no further answers":
+F = do(move(p2,’C’),do(move(p1,’D’),s0));
+F = do(move(p1,’D’),do(move(p2,’C’),s0));
+false.
+In the first solution, p1 acts first and p2 second, while in the second solution, the order is reversed.
+This outcome is expected, as both players have initial control in our game definition, resulting in both action order combinations being shown.
+3.5.3 Strategy
+Strategies are submitted separately from the game-dependent component.
+The select/4 predicate is used to determine the next move M of player P against opponent O in game situation S, as illustrated in the tit-for-tat strategy example below:
+select(P, O, S, M) if
+not holds(last_move(O, _LMo), S) and
+holds(default_move(P, M), S).
+select(P, O, S, Mo) if
+holds(last_move(O, Mo), S).
+Here, a player P selects the default move first (typically to cooperate), otherwise it mirrors the opponent’s move in the last round.
+holds/2 is based on the implementation of the Situation Calculus [22].
+The strategy relies on specific predicates—default_move/2 and last_move/2—which are required from the game-dependent part of the solver.
+More sophisticated strategies can be expressed.
+3.5.4 Constraints validation
+We check that the autoformalized payoff matrix satisfies the structural and content constraints of the specified game (as outlined in Sect.
+2.1), by defining validity constraints.
+The listing below specifies the payoff matrix validity constraints for the Prisoner’s Dilemma.
+valid_pd_payoffs(T,R,P,S,C,D) if
+payoff(C,C,R,R) and
+payoff(C,D,S,T) and
+T>R and
+payoff(D,C,T,S) and
+payoff(D,D,P,P) and
+R>P and
+P>S.
+Since the autoformalization module does not impose constraints on the names of actions, this predicate also identifies which actions names correspond to specific actions.
 4 Methods
 4.1 Parameters
+Table 2: Common experimental parameters.
 Parameter Value
@@
 Maximum attempts number $5$
-Table 2: Common experimental parameters.
+Table 3: Experiment-specific parameters. C stands for a mode in which an agent plays against its clone with an opposite strategy, while RR stands for round robin.
 Parameter Exp. 1 Exp. 2 Exp. 3 Exp. 4
@@
 Rounds $4$ $4$ $10$ $10$
-match maker S S RR S
-Table 3: Experiment-specific parameters. S stands for a mode in which an agent plays against its clone with an opposite strategy, while RR stands for round robin.
+Match maker C C RR C
 To evaluate the framework, we conducted four experiments: i) autoformalization of game scenarios with numerical (Exp.
@@
 Table 2 provides the experimental parameters shared across all experiments, while Table 3 details the parameters unique to each specific experiment.
+4.2 Dataset
+We used a dataset of 110 natural language game-theoretic scenarios, with and without numerical payoffs, based on an improved version of the dataset introduced in [23] (available in the repository).
+For both variants, the dataset included 5 scenarios per game employing common metaphors (e.g., two prisoners in the Prisoner’s Dilemma) for the five canonical games described in Section 2.1.
+The remaining 50 scenarios—10 per game—used alternative descriptions that deliberately avoided these typical metaphors (e.g., two politicians competing in a campaign rather than two prisoners in the Prisoner’s Dilemma).
 4.3 Autoformalization of Interaction Descriptions (Numerical Payoffs)
-We used a dataset of 55 natural language game-theoretic scenarios with numerical payoffs to evaluate the syntactic and semantic accuracy of generated logic programs.
-The dataset included 5 scenarios that employed common metaphors (e.g., two prisoners in the Prisoner’s Dilemma) for the five canonical games introduced in Section 2.1.
-The remaining 50 scenarios—10 for each game—used alternative descriptions that avoided these typical metaphors (e.g., two politicians competing in a campaign rather than two prisoners in the Prisoner’s Dilemma).
-For each scenario, five agents autoformalized game rules and then played tournament against their clones, generating all four possible pairs of moves in four rounds.
-Each agent had a maximum of five attempts to produce syntactically correct code; agents failing after five attempts were labelled as syntactically incorrect.
-Post-tournament, each agent’s individual and total payoffs were compared to the target, with matches indicating semantic correctness.
+For each scenario, five agents autoformalized the game rules and subsequently played tournaments against clones of themselves, generating all four possible pairs of moves across four rounds.
+Each agent was allowed up to five attempts to produce syntactically correct code; agents that failed after five attempts were labeled as syntactically incorrect.
+Following the tournament, each agent’s individual and total payoffs were compared to the target values, with exact matches indicating semantic correctness.
 4.4 Autoformalization of Interaction Descriptions (Non-Numerical Payoffs)
-This experiment was analogous to the previous one, but it utilized interaction scenarios without numerical payoffs (e.g., ”If both remain silent, they get minor sentences,” to describe the outcome of mutual cooperation in PD).
-Therefore, the autoformalizing LLM had to infer the appropriate payoffs correctly.
-Since it was not possible to provide a ground truth matrix in this case, we validated the semantic correctness by performing constraint checks (described in Sect.
-3.2) to determine whether the payoff matrix satisfied the constraints for the ground truth game type (Sect.
-2.1).
+This experiment followed the same procedure as the numerical payoffs evaluation but used interaction scenarios without explicit numerical payoffs (e.g., “If both remain silent, they receive minor sentences” to describe mutual cooperation in the Prisoner’s Dilemma).
+In these cases, the autoformalizing LLM was required to infer appropriate payoffs from the textual descriptions.
+As a ground truth payoff matrix could not be provided, semantic correctness was instead assessed through constraint checking (as described in Section 3.5), verifying whether the generated payoff matrix satisfied the properties required for the corresponding game type (Section 2.1).
 4.5 Axelrod’s Tournament
+Table 4: Strategies utilized in the evaluation.
 Strategy Description
@@
 tit-for-tat Start with a default move. Then, mirror the opponent’s move in the previous round.
-Table 4: Strategies utilized in the evaluation.
-To demonstrate the framework’s usefulness in analysing interaction scenarios, we selected five agents created during the first experiment, one for each of the five games.
-For each game, we conducted a tournament modelled after Axelrod’s tournament Axelrod and Hamilton (1981).
+To demonstrate the framework’s applicability for analysing interaction scenarios, we selected five agents generated during the first experiment, one representing each of the five canonical games.
+For each game, we conducted a tournament inspired by Axelrod’s seminal tournament [4].
 In this setup, six copies of each agent were created and assigned one of six strategies.
-Each agent then played against every other agent, including itself.
-The strategies used are listed in Table 4.
-The main metric for evaluating strategy performance was the total payoff achieved by an agent using a given strategy in each game.
-To allow for cross-game comparisons of strategy effectiveness, total payoffs were normalized using:
-$norm\_payoff=\frac{total\_payoff-min\_total\_payoff}{{max\_total\_payoff-min\_total\_payoff}},$
-$Total\_payoff$ refers to the sum of the payoffs an agent collects across all rounds and $min\_total\_payoff$ and $max\_total\_payoff$ represent the minimum and maximum total payoffs achievable, respectively.
+Each agent competed against every other agent, including itself.
+The strategies are listed in Table 4.
+The primary metric for evaluating strategy performance was the total payoff achieved by an agent using a particular strategy within each game.
+To enable cross-game comparisons of strategy effectiveness, total payoffs were normalized using the following formula:
+$\mathrm{NormPayoff}=\frac{\mathrm{TotalPayoff}-\mathrm{MinTotalPayoff}}{\mathrm{MaxTotalPayoff}-\mathrm{MinTotalPayoff}}$
+where TotalPayoff denotes the sum of payoffs accumulated by an agent across all rounds, and MinTotalPayoff and MaxTotalPayoff represent the minimum and maximum total payoffs achievable in that game, respectively.
 4.6 Autoformalization of Strategies
-We then evaluated the semantic correctness of the autoformalized strategies.
-To achieve this, we used the natural language description of the tit-for-tat strategy, along with its Prolog implementation, as an example.
-For the other five strategies, the autoformalization module was provided with their natural language descriptions, similar to those shown in Table 4.
-Each autoformalized strategy was then assigned to an agent with predefined Prisoner’s Dilemma game rules.
-This agent played four rounds against a clone utilizing the anti-tit-for-tat strategy.
-The agent’s total payoff was compared to a target payoff specific to each strategy to assess correctness.
-For the random strategy, correctness was also manually validated.
-This tournament setup was repeated five times.
+For the autoformalization of strategies, we used the natural language description of the tit-for-tat strategy, along with its Prolog implementation, as an example for one-shot learning.
+The autoformalization module then received natural language descriptions of the remaining five strategies, similar to those shown in Table 4, to generate their representations.
+Each autoformalized strategy was assigned to an agent equipped with predefined Prisoner’s Dilemma rules.
+The agent then played four rounds against a clone using the anti-tit-for-tat strategy.
+As in the game autoformalization experiments, the agent’s total payoff was compared against a target payoff specific to each strategy to assess semantic correctness.
+For the random strategy, correctness was additionally verified through manual inspection.
 5 Results and Discussion
 5.1 Autoformalization of Interaction Descriptions
-1 2 3 4 5
+Table 5: Percentage of number of attempts at creating syntactically correct code for each model.
+1 (%) 2 (%) 3 (%) 4 (%) 5 (%)
 Numeric payoffs
-Claude 3.5 S 100.00% 0.00% 0.00% 0.00% 0.00%
-GPT-4o 96.77% 2.87% 0.00% 0.00% 0.36%
+Claude 3.5 S $100.00$ $0.00$ $0.00$ $0.00$ $0.00$
+GPT-4o $96.77$ $2.87$ $0.00$ $0.00$ $0.36$
 Non-numeric payoffs
-Claude 3.5 S 100.00% 0.00% 0.00% 0.00% 0.00%
-GPT-4o 97.64% 2.02% 0.34% 0.00% 0.00%
-Table 5: Distribution of attempts at creating syntactically correct code for each model.
-Table 5 shows the distribution of attempts needed to generate syntactically correct code for both models.
-The type of description (numeric vs.
-non-numeric) had no influence on syntactic correctness of the generated code.
-Claude 3.5 Sonnet achieved syntactically correct code in 100% of cases and GPT-4o in nearly 100%.
-The few errors observed involved non-Prolog comment characters and non-escaped special characters in strings, all of which were successfully corrected on the second attempt.
-This demonstrates that providing a specific line that generated the error in the feedback prompt significantly enhances an LLM’s ability to self-correct compared to merely providing an error message Mensfelt et al.
-(2024).
-The only case where GPT-4o was unable to correct a syntactic error involved the error message “Unexpected end of file“ which indicated the end of the file as the error’s source.
-However, the actual line generating the error was not provided in the feedback prompt.
-Syntactic Runtime Semantic
+Claude 3.5 S $100.00$ $0.00$ $0.00$ $0.00$ $0.00$
+GPT-4o $97.64$ $2.02$ $0.34$ $0.00$ $0.00$
+Table 5 presents the distribution of attempts required to generate syntactically correct code for both models.
+Claude 3.5 Sonnet achieved syntactically correct code on the first attempt in 100% of cases, while GPT-4o did so in 97% of cases.
+The few errors observed in GPT-4o primarily involved non-Prolog comment characters or unescaped special characters in strings, most of which were corrected on the second attempt.
+These results demonstrate that providing the specific Prolog line that triggered the error in the feedback prompt significantly improves the LLM’s ability to self-correct, compared to providing only an error message [23].
+The only instance where GPT-4o failed to correct a syntactic error involved an “Unexpected end of file” message.
+In this case, the feedback prompt did not include the problematic line itself, but only the generic error message, limiting the model’s ability to identify and fix the issue.
+Table 6: Syntactic, runtime, and semantic correctness for both models and types of scenario descriptions.
+Syntactic (%) Runtime (%) Semantic (%)
 Numeric payoffs
-Claude 3.5 Sonnet 100.0% 88.0% 86.0%
-GPT-4o 100.0% 90.0% 86.0%
+Claude 3.5 Sonnet 100.0 88.0 86.0
+GPT-4o 100.0 90.0 86.0
 Non-numeric payoffs
-Claude 3.5 Sonnet 100.0% 79.0% 67.0%
-GPT-4o 100.0% 91.0% 68.0%
-Table 6: Syntactic, runtime, and semantic correctness for both models and types of scenario descriptions.
+Claude 3.5 Sonnet 100.0 79.0 67.0
+GPT-4o 100.0 91.0 68.0
 The type of description had no influence on runtime errors in GPT-4o (runtime correctness of 90% and 91% for numeric and non-numeric payoffs, respectively, Table 6).
@@
 This highlights the need to extend the autoformalization feedback loop to the runtime stage and leverage the tracing mechanism of the solver to identify and create missing predicates.
-Numeric Payoffs
-BS HD MP PD SH
-Syntactic 100.0% 100.0% 100.0% 100.0% 100.0%
-Runtime 89.1% 96.4% 68.5% 96.4% 87.3%
-Semantic 89.1% 96.4% 68.5% 89.1% 87.3%
-Non-Numeric Payoffs
-BS HD MP PD SH
-Syntactic 100.0% 100.0% 100.0% 100.0% 100.0%
-Runtime 82.1% 85.7% 74.0% 70.7% 83.1%
-Semantic 82.1% 77.8% 34.2% 64.0% 83.1%
 Table 7: Correctness by game for Claude 3.5 Sonnet for game descriptions with numeric and non-numeric payoffs.
 Numeric Payoffs
-BS HD MP PD SH
-Syntactic 98.3% 100.0% 100.0% 100.0% 100.0%
-Runtime 83.1% 89.1% 96.4% 89.1% 90.9%
-Semantic 79.7% 89.1% 90.9% 85.5% 83.6%
+BS (%) HD (%) MP (%) PD (%) SH (%)
+Syntactic $100.0$ $100.0$ $100.0$ $100.0$ $100.0$
+Runtime $89.1$ $96.4$ $68.5$ $96.4$ $87.3$
+Semantic $89.1$ $96.4$ $68.5$ $89.1$ $87.3$
 Non-Numeric Payoffs
-BS HD MP PD SH
-Syntactic 100.0% 100.0% 100.0% 100.0% 100.0%
-Runtime 93.2% 88.1% 87.3% 96.5% 91.5%
-Semantic 93.2% 67.8% 30.2% 73.7% 76.3%
+BS (%) HD (%) MP (%) PD (%) SH (%)
+Syntactic $100.0$ $100.0$ $100.0$ $100.0$ $100.0$
+Runtime $82.1$ $85.7$ $74.0$ $70.7$ $83.1$
+Semantic $82.1$ $77.8$ $34.2$ $64.0$ $83.1$
 Table 8: Correctness by game for GPT-4o for game descriptions with numeric and non-numeric payoffs.
-The semantic correctness was lower than runtime correctness for both models and description types (Table 6).
-This means that even though an agent was able to select a move during runtime and update its state, the semantics of the program were not correct.
-For descriptions with numerical payoffs, this implies that the payoff matrix does not contain the required payoffs.
-For descriptions without numerical payoffs, it indicates that the payoff matrix does not satisfy the constraints of the required game.
-Tables 7 and 8 present the syntactic, runtime, and semantic correctness for different games.
-The runtime and semantic correctness varied depending on the model and the game.
-Notably, Matching Pennies consistently resulted in low semantic correctness for non-numerical descriptions, with 34.2% correctness for Claude 3.5 Sonnet and 30.2% correctness for GPT-4o.
-This may be attributed to some descriptions specifying which player wins if the choices match or mismatch, without explicitly stating that the other player will lose—a constraint of the game.
-Extending the autoformalization feedback loop to include semantic evaluation has the potential to significantly improve correctness.
+Numeric Payoffs
+BS (%) HD (%) MP (%) PD (%) SH (%)
+Syntactic $98.3$ $100.0$ $100.0$ $100.0$ $100.0$
+Runtime $83.1$ $89.1$ $96.4$ $89.1$ $90.9$
+Semantic $79.7$ $89.1$ $90.9$ $85.5$ $83.6$
+Non-Numeric Payoffs (%)
+BS (%) HD (%) MP (%) PD (%) SH (%)
+Syntactic $100.0$ $100.0$ $100.0$ $100.0$ $100.0$
+Runtime $93.2$ $88.1$ $87.3$ $96.5$ $91.5$
+Semantic $93.2$ $67.8$ $30.2$ $73.7$ $76.3$
+Semantic correctness was lower than runtime correctness for both models and description types (Table 6).
+This indicates that, although an agent was able to select moves and update its state during execution, the underlying semantics of the program were incorrect.
+For descriptions with numerical payoffs, semantic errors imply that the generated payoff matrix did not contain the required values.
+For descriptions without numerical payoffs, they indicate that the payoff matrix failed to satisfy the constraints associated with the intended game.
+Detailed syntactic, runtime, and semantic correctness results for different games are shown in Tables 7 and 8.
+Runtime and semantic correctness varied across models and games.
+Notably, Matching Pennies consistently exhibited low semantic correctness for non-numerical descriptions, with only 34.2% correctness for Claude 3.5 Sonnet and 30.2% for GPT-4o.
+This may be attributed to certain descriptions specifying only which player wins when choices match or mismatch, without explicitly stating that the other player loses—a crucial constraint of the game.
+Incorporating semantic evaluation into the autoformalization feedback loop may substantially improve correctness by addressing such issues.
 5.2 Axelrod’s Tournament
 Figure 2: Normalized total payoff for each strategy by game.
+Rows are ordered by the strategies with the highest average total payoff, and columns are arranged similarly based on the games.
+The total payoff for the winning strategy in each game is highlighted in white.
 Figure 2 presents the normalized total payoffs for each strategy across all games.
-Overall, the best-response strategy proved to be the most effective on average, followed closely by tit-for-tat.
-Best-response was most effective in all games except the Battle of the Sexes and the Prisoner’s Dilemma.
-In the Prisoner’s Dilemma, best-response achieved a lower payoff than the anti-default-move strategy (analogous to “defect”), as it initiated with the “cooperate” move.
-The results demonstrate that agents with autoformalized rules can effectively simulate and assess the performance of various strategies within different modelled scenarios.
+Overall, the best-response strategy proved the most effective on average, followed closely by tit-for-tat.
+Best-response achieved the highest scores in Stag Hunt and Matching Pennies (tied with tit-for-tat in the latter).
+However, in the Prisoner’s Dilemma, it earned a lower payoff than the anti-default-move strategy (analogous to “Defect”), as it opened with a “Cooperate” move.
+These results demonstrate that agents equipped with autoformalized rules can successfully simulate strategic interactions and assess the relative performance of various strategies across different modelled scenarios.
 5.3 Autoformalization of Strategies
-Strategy Claude 3.5 Sonnet GPT-4o
-anti-default-move 100.00% 100.00%
-anti-tit-for-tat 100.00% 100.00%
-best-response 80.00% 100.00%
-default-move 100.00% 100.00%
-random 80.00% 100.00%
-Table 9: Percentage of semantic correctness for each strategy.
+Table 9: Percentage of semantic correctness for each autoformalized strategy.
+Strategy Claude 3.5 Sonnet (%) GPT-4o (%)
+anti-default-move $100.00$ $100.00$
+anti-tit-for-tat $100.00$ $100.00$
+best-response $80.00$ $100.00$
+default-move $100.00$ $100.00$
+random $80.00$ $100.00$
 Table 9 presents the percentage of correctly autoformalized strategies.
-GPT-4o achieved 100% correctness in autoformalizing all five strategies.
-Claude 3.5 Sonnet produced one incorrect strategy for best-response and one for random, which are the strategies that deviate the most from the tit-for-tat strategy presented as an example in the prompt.
-One error occurred during runtime, and another during semantic validation.
-Overall, the high correctness demonstrates that the GAMA framework has the potential to effectively autoformalize both game rules and gameplay strategies.
-6 Related Work
-Originally proposed for agents in game scenarios, the Game Description Language (GDL) offers a promising approach for formalising complex strategic interactions in multi-agent environments by providing a declarative framework to define game rules, legal moves, and outcome conditions Genesereth et al.
-(2005).
-Agents that employ GDL are often used to systematically explore game states and strategic options, supporting rigorous analysis across both cooperative and competitive games, as suggested by previous approaches (e.g., see  Schiffel and Thielscher (2009)).
-However, these approaches typically rely on human-driven formalisation of specific games, a limitation addressed by the LLM-augmented agents and the GDL representation used here, which enable autoformalization and enhance flexibility.
-There is a growing body of research exploring the use of LLMs in game theory Lorè and Heydari (2024), systematically examining their potential as player agents in game-theoretic scenarios Fan et al.
-(2024), utilising their inference abilities to make decisions in games with imperfect information Guo et al.
-(2023), comparing their behaviour with human participants Fontana et al.
-(2024), and deploying them in simulation frameworks Mao et al.
-(2024).
-While we share with these works an interest in the broad capabilities and value of LLMs in game theoretic settings, our focus diverges by concentrating on how to use LLMs to generate GDL-like executable game specifications.
-These formal specifications enable agents to rigorously analyse game situations and reason about possible outcomes, providing a structured and robust foundation for decision-making within game-theoretic frameworks.
-Our work builds on previous research discussed in Mensfelt et al.
-(2024), significantly extending it by embedding the autoformalization module within an agent model, incorporating the autoformalization of player strategies, implementing a distinct correctness validation process for the autoformalization through tournament play, and conducting experiments to validate the results.
-Autoformalization Wu et al.
-(2022), has shown promising results in translating natural language into mathematical formalisms, enabling the integration of large language models with formal tools for solving reasoning tasks Wu et al.
-(2022); He-Yueya et al.
-(2023); Jiang et al.
-(2022).
-This approach has also been successfully extended to creating logical representations from natural language descriptions.
-For instance, Yang et al.
-(2023) applied GPT-3 with few-shot prompting to translate natural language sentences into formal representations, which were then inputted into answer set programs.
-While effective on NLP benchmarks, this method faced occasional translation errors.
-To further enhance performance, Pan et al.
-(2023) combined in-context learning with self-refinement based on solver evaluations, surpassing standard LLM and chain-of-thought prompting approaches on logical reasoning benchmarks.
-In the context of temporal logic, Cosler et al.
-(2023) addressed the problem of translation errors with an interactive system allowing users to refine sub-translations.
-Fine-tuning LLMs has also proven effective for boosting accuracy in translations to temporal logics Chen et al.
-(2023) and deductive datasets Feng et al.
-(2023).
+GPT-4o achieved 100% correctness across all five strategies.
+Claude 3.5 Sonnet produced one incorrect autoformalization each for best-response and random, which are the most divergent from the tit-for-tat strategy provided as an example in the prompt.
+Of these errors, one occurred during runtime and the other during semantic validation.
+Overall, the high levels of correctness demonstrate that the GAMA framework can effectively autoformalize both game rules and gameplay strategies, including those with varying levels of similarity to the examples provided.
 6 Conclusions and Future Work
-We have presented a framework for the autoformalization of interaction scenarios, specifically those modelled by $2\times 2$ simultaneous-move games.
-The framework enables the autoformalization of strategies for these games and the simulation of tournaments based on the autoformalized rules.
-The generated programs undergo three levels of validation: syntactic validation by consulting a solver, functional testing through gameplay, and semantic validation by comparing achieved payoffs with target outcomes or by checking the game constraints.
-Evaluation of the framework demonstrated high syntactic and semantic accuracy in the code generated by agents across various interaction scenarios.
-Future work will focus on extending the framework to model games beyond $2\times 2$ simultaneous-move scenarios.
-Our previous work Mensfelt et al.
-(2024) demonstrated that using our sample solver for the Prisoner’s Dilemma we could successfully generate code for structurally similar games, such as rock-paper-scissors (with a different number of moves) and sequential versions of the PD.
-However, achieving successful generalization to a broader class of games will require assembling a more diverse set of in-context learning examples and developing methods to retrieve the most relevant examples for autoformalization in specific scenarios.
-Another direction for future work involves extending the autoformalization feedback loop to include runtime and semantic validation checks, further improving the correctness of the generated programs.
-These extensions will provide a robust framework for generating reasoning modules for decision-making agents in a wider class of games and game-like interactions.
-Acknowledgments
+We have proposed a framework for the autoformalization of interaction scenarios, focusing on those modelled as $2\times 2$ simultaneous-move games.
+The framework supports autoformalization of both game rules and strategies, and enables simulation of tournaments based on the generated formal representations.
+Programs produced by the framework undergo three levels of validation: syntactic validation through solver consultation, runtime testing via gameplay, and semantic validation based on target payoffs or game-specific constraints.
+Experimental results demonstrated high syntactic and semantic accuracy across diverse interaction scenarios.
+Future work will extend the framework beyond $2\times 2$ simultaneous-move games.
+Our previous work [23] showed that generalization to structurally similar games, such as rock-paper-scissors and sequential versions of the Prisoner’s Dilemma, could be addressed using the same solver.
+Generalization to a broader class of games, however, will require assembling a more diverse collection of in-context learning examples and developing retrieval mechanisms to select relevant examples for autoformalization in specific scenarios.
+Another promising direction involves expanding the feedback loop to incorporate runtime and semantic validation, further enhancing the correctness and robustness of the generated programs.
+These extensions aim to establish a more general and reliable framework for generating reasoning modules, enabling decision-making agents to operate effectively across a wider range of games and interaction scenarios.
+{ack}
 This work was supported by a Leverhulme Trust International Professorship Grant (LIP-2022-001).